### v1

# Bayesian Empirical Bayes: Simultaneous Inference from Probabilistic Symmetries

Abstract
Empirical Bayes (EB) improves the accuracy of simultaneous inference “by learning from the experience of others” (Efron2012). Classical EB theory focuses on latent variables that are iid draws from a fitted prior (Efron2019OracleBayes). Modern applications, however, feature complex structure, like arrays, spatial processes, or covariates. How can we apply EB ideas to these settings? We propose a generalized approach to empirical Bayes based on the notion of probabilistic symmetry. Our method pairs a simultaneous inference problem–with an unknown prior–to a symmetry assumption on the joint distribution of the latent variables. Each symmetry implies an ergodic decomposition, which we use to derive a corresponding empirical Bayes method. We call this method Bayesian empirical Bayes (BEB). We show how BEB recovers the classical methods of empirical Bayes, which implicitly assume exchangeability. We then use it to extend EB to other probabilistic symmetries: (i) EB matrix recovery for arrays and graphs; (ii) covariate–assisted EB for conditional data; (iii) EB spatial regression under shift invariance. We develop scalable algorithms based on variational inference and neural networks. In simulations, BEB outperforms existing approaches to denoising arrays and spatial data. On real data, we demonstrate BEB by denoising a cancer gene–expression matrix and analyzing spatial air-quality data from New York City.
11footnotetext: Department of Statistics, Columbia University, New York, NY22footnotetext: Department of Chemistry, Technical University of Denmark, Denmark33footnotetext: Department of Computer Science, Columbia University, New York, NY
Contact: bw2766@columbia.edu, ew2760@columbia.edu, david.blei@columbia.edu
1 Introduction
Empirical Bayes (EB) is about simultaneous posterior inference. We observe related data $x_{1},\ldots,x_{n}$, where each data point is associated with a latent variable $z_{i}$, and assumed drawn from $x_{i}\sim\mathrm{p}(x\,|\,z_{i})$. The form of the conditional likelihood is known. The goal of EB is to use Bayesian inference to estimate all the $z_{i}$ together (Robbins1956; Efron2015FreqBayes; Efron2019OracleBayes).
In a traditional Bayesian analysis, we specify a prior $z_{i}\sim\textrm{g}(z)$ and compute the posteriors $\mathrm{p}(z_{i}\,|\,x_{i})$. But in EB, we treat the prior $\textrm{g}(z)$ as unknown and we estimate it from the data. EB thus blends Bayesian and frequentist thinking: we reason in a frequentist way about the prior in order to form a collection of Bayesian estimates.
[…]
The assumption $z_{1},\ldots,z_{n}\stackrel{{\scriptstyle iid}}{{\sim}}g$ is central to EB thinking. It says that the latent variables share a common prior, which justifies the EB procedure. But many modern datasets violate this iid assumption. Latent variables may come in rows and columns, they may lie on a spatial or temporal grid, or they may depend on known covariates. While these settings clearly violate the iid assumption, they still exhibit symmetries that tie the latent variables together. In this paper, we develop a general way to use such probabilistic symmetries to extend EB to non-iid settings.
Stepping back, the assumption that $z_{1},\ldots,z_{n}$ are iid from g is a special case of a broader principle: exchangeability. If the joint distribution of the latent variables is invariant to permutations, then—under mild conditions—it can be viewed as part of an infinitely exchangeable sequence. The iid structure then arises through de Finetti’s famous representation theorem (definetti1929). It states that any exchangeable sequence can be written as a mixture of iid laws,
$\mathrm{p}(z_{1},\ldots,z_{n})=\int\Bigl[\prod_{i=1}^{n}\mathrm{p}(z_{i}\mid\textrm{g})\Bigr]\,\mathrm{d}\mu(\textrm{g}),$ (1)
where g is a probability distribution on $\mathcal{Z}$ and $\mu$ is a distribution over such distributions. De Finetti’s result is important because it connects a symmetry assumption—namely, exchangeability—to classical Bayesian reasoning (Savage1954; Robert2007Bayesianchoice). If the $z_{i}$’s are exchangeable, then the prior $g$ appears as the latent variable that renders the $z_{i}$ conditionally independent.
Our insight is that empirical Bayes uses the same hierarchical structure as in Eq.˜1, where $z_{i}$ are conditionally iid given $g$. However, in place of integration with respect to $\mu$, EB substitutes a fitted value $\hat{\textrm{g}}$ obtained from data. Seen through de Finetti’s theorem, we can view EB as a Bayesian inference under exchangeability, but with the mixing distribution collapsed to a point estimate.
Using this perspective, we extend classical EB by relaxing its exchangeability assumption to other probabilistic symmetries. Latent variables often have structure, such as being organized in an array or indexed by space or time. Such settings are not fully exchangeable, but they may satisfy weaker invariances. An array may be invariant to permutations of its rows and columns; a spatial process may be invariant to shifts in location (Kallenberg2005book). Each symmetry has an analogue of de Finetti’s representation theorem. We will use these representation theorems to extend EB to non-iid settings.
We call our approach Bayesian empirical Bayes (BEB). We begin with a dataset $\mathbf{x}$ and an assumed probabilistic symmetry. We posit that each observation arises from a known likelihood conditioned on a local latent variable $\mathrm{p}(x_{i}\,|\,z_{i})$, and our goal is to infer all $z_{i}$ simultaneously. The key idea is that our assumed symmetry implies an ergodic decomposition of the prior, i.e., a generalization of de Finetti’s theorem (Orbanz2015). Specifically, if the latent variables are invariant under a group $\Phi$ acting on its index set, then their joint distribution admits the form
$\mathrm{p}(\bm{z})=\int\mathrm{p}(\bm{z}\mid\textrm{g})\,\mathrm{d}\mu(\textrm{g}),$ (2)
where g indexes the $\Phi$-ergodic family. The chosen symmetry determines the form of g and the conditional distribution $\mathrm{p}(\bm{z}\,|\,\textrm{g})$.
Following the EB recipe, we fit $g$ by maximum marginal likelihood under the expanded model $\mathrm{p}(\bm{z},\bm{x})$,
$\hat{\textrm{g}}\in\operatorname*{arg\,max}_{\textrm{g}}\ \log\int\mathrm{p}(\bm{x}\,|\,\bm{z})\,\mathrm{p}(\bm{z}\,|\,\textrm{g})\,\mathrm{d}\bm{z}.$ (3)
[…]
We illustrate BEB in several data analysis settings. For array data, the relevant symmetry is separate or joint exchangeability, and the ergodic decomposition is given by the Aldous-Hoover theorem (Aldous1985). The invariant variable g is a measurable function on $0,1]^{3}$, which we parameterize with a neural network. For spatial data, the relevant symmetry is shift invariance, and we can flexibly model g with stationary ergodic Gaussian processes. In both settings, we can additionally condition on covariates and derive covariate-dependent EB procedures.
Thus, Bayesian empirical Bayes produces a generalization of EB to any probabilistic symmetry. The invariant variable g replaces the classical prior, and the ergodic decomposition supplies the structure needed for inference.
The rest of the paper is organized as follows. §˜2 introduces the modeling assumptions and the general methodology of Bayesian empirical Bayes. §˜3 develops new inference methods from Bayesian EB principles in three settings: the classical normal sequence model, empirical Bayes matrix recovery, and empirical Bayes spatial regression. For normal means, we show that Bayesian EB with an exchangeable sequence model recovers the nonparametric maximum-likelihood method of Efron2015deconvolution. For the latter two, we derive new EB procedures based on alternative invariances-joint/separate/relative exchangeability for matrices and stationarity for spatio-temporal data. §˜4 evaluates these methods in simulation studies. §˜5 applies them to real datasets. §˜6 concludes with discussion and future directions. Detailed algorithms for our inference procedures and all technical proofs appear in the Appendix.
1.1 Related Work
Empirical Bayes (EB) dates back to the seminal works of Robbins1951compound; Robbins1956 and has been developed in both parametric and nonparametric forms (Efron1971; Efron1976; Efron2001; Brown2009; Jiang2009; Efron2011; Efron2012; Efron2014; Efron2015deconvolution; Efron2021; Soloff2021; Ignatiadis2022; Yang2024). Much of the nonparametric theory centers on normal and Poisson sequence models. For normal means, Efron2015FreqBayes surveys frequentist assessments of EB estimators and Efron2019OracleBayes surveys oracle–Bayes viewpoints. Beyond g-modeling (modeling the prior), Efron2014 described an alternative EB approach via $f$-modeling, which models the marginal data distribution; a key tool is Tweedie’s formula, linking the posterior mean to the derivative of the log marginal density. For the $f$-modeling methods, Saremi2019 and ghosh2025stein obtain nonparametric EB estimators using flexible score functions, by minimizing the score-matching loss. EB also underpins large–scale multiple testing (Efron2012; Stephens2017; Ignatiadis2021). While it is impossible to exhaust the works on empirical Bayes for the normal sequence model, we refer to Ignatiadis2025EB for a recent textbook overview. In this work we focus on g-modeling and extend it beyond the iid prior setting to accommodate general invariant latent structures.
Nonparametric EB methods for non-sequence models have only appeared more recently. In high–dimensional linear models and variable selection, variational EB methods—such as SuSiE and related procedures—learn a nonparametric prior over coefficients (Carbonetto2012; Wang2020susie; Kim2022). The prior is exchangeable with respect to covariates (trippe2021high). The theoretical properties of these methods are beginning to be understood (Mukherjee2023; Fan2023). For low–rank matrices, empirical Bayes matrix factorization and EB–PCA fit learned shrinkage priors for loadings and scores via approximate inference (Wang2021EBMF; Zhong2022; denault2025covariate). These new developments share the theme of fitting shared priors for latent variables to structure the borrowing of information. Inspired by this line of work, we provide a unified framework for inferring many latent variables by learning a shared latent structure defined by probabilistic symmetries. This invariance-based perspective enables systematic extensions of EB methods.
From a Bayesian perspective, EB concerns approximating a full Bayesian posterior by a point estimate of the hyperparameter (Ch. 5.6.4 of (Bernardo1994)) and is closely connected to hierarchical Bayes (Lindley1972; good1992introduction). Deely1981 compare the statistical performance of parametric empirical Bayes and hierarchical Bayes on simultaneous inference problems. One can place additional layers of priors on EB hyperparameters; for example, Kass1989EB introduce a hyperprior to capture uncertainty in EB estimation of prior hyperparameters in general hierarchical models. Rousseau2017 provide sufficient conditions for general consistency and posterior contraction rates of the EB posterior under parametric maximum marginal likelihood; in their setting, the results show that hierarchical and empirical Bayes posteriors obtain comparable posterior contraction rates. Rizzelli2024 analyze the asymptotic behavior of posteriors obtained after fitting such hyperparameters. Buta2011 and Doss2024 discussed the connection between empirical Bayes and Bayesian sensitivity analysis. While the Bayesian literature on EB has focused almost exclusively on parametric models, Bayesian empirical Bayes provides a principled approach to nonparametric empirical Bayes, motivated from Bayesian reasoning.
Many objects in classical statistics obey probabilistic symmetries (see Diaconis1988Book; see also Chapter 3 of Lehmann2006). The notion of exchangeability, for example, serves as the foundation for Bayesian theory (definetti1929; Savage1954; Bernardo1994) and underlies much of practical Bayesian modeling (Gelman1995). Other applications of group symmetries include invariant testing—e.g., permutation-based testing methods (eden1933; fisher1935; pitman1937; pesarin2001; ernst2004; pesarin2010; pesarin2012; good2006; anderson2001; kennedy1995; hemerik2018; ramdas2023). Early advances for estimation and testing under general groups include lehmannstein1949 and hoeffding1952; the celebrated Hunt-Stein theorem states conditions under which a minimax invariant test exists in hypothesis testing (Lehmann1959). Broader treatments appear in eaton1989; wijsman1990; giri1996. In modern probabilistic modeling, the ergodic decomposition theorem provides a foundation for building invariant representations and modeling (Orbanz2015). For example, Lloyd2012 study exchangeable arrays by placing a Gaussian process prior on the underlying random function. We build on the general theory of probabilistic symmetries as developed in Kallenberg2005book and Orbanz2015. The theory characterizes general invariant distributions under group actions using ideas from ergodic theory. It allows us to extend empirical Bayes to many different kinds of structured data settings, creating a foundation for new applications of EB.
2 Bayesian Empirical Bayes
Consider a structured dataset $\bm{x}=(x_{\omega})_{\omega\in\Omega}$, where each $x_{\omega}$ is indexed by an element in an index set $\Omega$. For example, if $\Omega=\mathbb{N}$, then $\bm{x}$ is an infinite sequence; if $\Omega=\mathbb{N}^{2}$, then $\bm{x}$ is an infinite array. (After describing a model of infinite data, we will reduce it to one of finite data.)
We posit that the data $\bm{x}$ are noisy measurements of a matching set of structured latent variables $\bm{z}=(z_{\omega})_{\omega\in\Omega}$. In the likelihood, each $x_{\omega}$ is conditionally independent given its local latent $z_{\omega}$,
$\mathrm{p}(\bm{x}\,|\,\bm{z})=\prod_{\omega\in\Omega}\mathrm{p}(x_{\omega}\,|\,z_{\omega}).$ (4)
We assume the form of the noise distribution $\mathrm{p}(x_{\omega}\,|\,z_{\omega})$ is known.
We place a prior distribution on the latent variables $\bm{z}\sim\mathrm{p}(\bm{z})$. The joint distribution of $\bm{z}$ and $\bm{x}$ is now
$\mathrm{p}(\bm{z},\bm{x})=\mathrm{p}(\bm{z})\prod_{\omega\in\Omega}\mathrm{p}(x_{\omega}\,|\,z_{\omega}).$ (5)
Note, we do not assume independence in the prior.
Our goal is simultaneous probabilistic inference of all local latent variables across $\Omega$. That is, we want to calculate the posterior $\mathrm{p}(\bm{z}\,|\,\bm{x})$. From this posterior, we can calculate summary statistics, such as the posterior means $\mathbb{E}\left[\bm{z}\,|\,\bm{x}\right]$, or other quantities of interest, such as credible intervals, quantiles, posterior predictive distributions, or event probabilities. In short, we can denoise the data.
We now make the empirical Bayes assumption: there exists a prior $\mathrm{p}^{\star}(\bm{z})$ such that the true data-generating distribution of $\bm{x}$ is expressed by marginalizing over $\bm{z}$.
Assumption 1 (Empirical Bayes).
There exists a prior $\mathrm{p}^{\star}(\bm{z})$ such that the true data-generating distribution can be written as
$\mathrm{p}^{\star}(\bm{x})=\int_{\mathcal{X}^{\Omega}}\mathrm{p}^{\star}(\bm{z})\prod_{\omega\in\Omega}\mathrm{p}(x_{\omega}\,|\,z_{\omega})\,\mathrm{d}\bm{z}.$
Of course, we do not know this true prior ahead of time - we just assume it exists. We call this the empirical Bayes assumption because it links the Bayesian model to the true data-generating process.
2.1 Priors with Probabilistic Symmetries
How do we set the prior $\mathrm{p}(\bm{z})$? We will use the idea of a probabilistic symmetry. A symmetry describes how we can relabel or transform the indices of structured random variables without changing their joint distribution.
Formally, let $\Phi$ denote a group of transformations acting bijectively on the index set $\Omega$, where each element $\phi\in\Phi$ reindexes the coordinates of $\Omega$. The action of $\phi$ on $\bm{z}$ is the index transformation
$(\phi(\bm{z}))_{\omega}=z_{\phi^{-1}(\omega)},\quad\omega\in\Omega.$ (6)
As examples, if $\Omega=\mathbb{N}$, then $\phi(\bm{z})$ reorders the elements of a sequence. If $\Omega=\mathbb{N}^{2}$, then $\phi(\bm{z})$ could permute rows and columns of a matrix. If $\Omega=\mathbb{R}^{2}$, then $\phi(\bm{z})$ could implement a spatial shift.
A distribution $\mathrm{p}(\bm{z})$ is $\Phi$-invariant if it does not change under these transformations. For example, an exchangeable sequence is invariant under permutations; a stationary process is invariant under spatial shifts.
Definition 1 ($\Phi$-invariance).
A joint distribution $\mathrm{p}(\bm{z})$ of structured variables $\bm{z}\in\mathcal{Z}^{\Omega}$ is $\Phi$-invariant if
$\phi(\bm{z})\overset{d}{=}\bm{z}\quad\text{for all }\phi\in\Phi,$ (7)
where equality is in distribution for all finite-dimensional marginals.
A formal consequence of $\Phi$-invariance is the ergodic decomposition theorem, which states that any $\Phi$-invariant distribution can be expressed as a mixture of ergodic distributions.
Theorem 1 (Ergodic decomposition; Varadarajan1963).
Let $\Phi$ be a nice group acting on $\mathcal{Z}^{\Omega}$. There exists a family of distributions
$\mathcal{E}:=\left\{\mathrm{p}(\bm{z}\,|\,\textrm{g}):\textrm{g}\in\mathcal{G}\right\}\subseteq\mathcal{P}(\mathcal{Z}^{\Omega})$ (8)
such that any $\Phi$-invariant distribution $\mathrm{p}(\bm{z})$ admits the unique decomposition
$\mathrm{p}(\bm{z})=\int_{\mathcal{G}}\mathrm{p}(\bm{z}\,|\,\textrm{g})\,\mathrm{d}\mu(\textrm{g}),$ (9)
for some distribution $\mu$ on $\mathcal{G}$. The set $\mathcal{E}$ is called the $\Phi$-ergodic family and its index g is called the invariant variable. [^1] It is also known as the directing random measure (Kallenberg2005book).
A “nice” group is a technical condition; see Appendix˜A for a formal definition. The $\Phi$-ergodic family is determined by the group $\Phi$ and forms the set of extreme points of the convex set of $\Phi$-invariant distributions on $\mathcal{Z}^{\Omega}$ (Kallenberg1997foundations, Theorem 25.24). Once $\Phi$ is fixed, so too is the form of $\mathrm{p}(\bm{z}\,|\,\textrm{g})$.
The ergodic decomposition theorem generalizes de Finetti’s representation (definetti1929) from exchangeable sequences to arbitrary group symmetries. Under exchangeability, the ergodic measures are precisely the i.i.d. product distributions.
The ergodic decomposition describes the structure of any $\Phi$-invariant distribution. We now apply this idea to the true prior.
Assumption 2 (True $\Phi$-invariant prior).
The true prior $\mathrm{p}^{\star}(\bm{z})$ is $\Phi$-invariant.
By Theorem˜1, there exists an invariant variable $\textrm{g}\in\mathcal{G}$ with distribution $\textrm{g}\sim\mu^{\star}$ such that $\mathrm{p}^{\star}(\bm{z})$ is a $\mu^{\star}$-mixture of the $\Phi$-ergodic family,
$\mathrm{p}^{\star}(\bm{z})=\int_{\mathcal{G}}\mathrm{p}(\bm{z}\,|\,\textrm{g})\,\mathrm{d}\mu^{\star}(\textrm{g}).$ (10)
Together, Assumptions˜1 and 2 state that among all models of the form in Eq.˜5, there exists a correctly specified version whose prior is $\Phi$-invariant. In the Bayesian empirical Bayes model, the latent variables $\bm{z}$ capture the structured, invariant part of the world, and the observations $\bm{x}$ are noisy realizations through the shared likelihood $\mathrm{p}(x_{\omega}\,|\,z_{\omega})$.
An important consequence of Assumption 2 is that the data-generating distribution $\mathrm{p}^{\star}(\bm{x})$ inherits the same symmetry as the prior. To see this, we first note that applying any transformation $\phi\in\Phi$ to both $\bm{x}$ and $\bm{z}$ leaves the joint distribution unchanged,
$\mathrm{p}^{\star}(\phi(\bm{x}),\phi(\bm{z}))$ $=\mathrm{p}^{\star}(\bm{z})\prod_{\omega\in\Omega}\mathrm{p}(x_{\phi^{-1}(\omega)}\,|\,z_{\phi^{-1}(\omega)})$ (11)
$=\mathrm{p}^{\star}(\bm{z})\prod_{\omega\in\Omega}\mathrm{p}(x_{\omega}\,|\,z_{\omega})=\mathrm{p}^{\star}(\bm{x},\bm{z}).$ (12)
Then we marginalize out $\bm{z}$ to give $\mathrm{p}^{\star}(\phi(\bm{x}))=\mathrm{p}^{\star}(\bm{x})$.
Alternatively, instead of assuming the true prior is invariant, we could assume that the marginal distribution of the data is invariant. Then, under mild regularity assumptions on the likelihood, the $\Phi$-invariance of the population distribution $\mathrm{p}^{\star}(\bm{x})$ also implies the $\Phi$-invariance of $\mathrm{p}^{\star}(\bm{z})$.
2.2 Examples of $\Phi$-invariant priors
We present three examples of $\Phi$-invariant priors, which we will use throughout this paper. In each, we specify the index set $\Omega$, the symmetry group $\Phi$, the form of the ergodic decomposition, and the corresponding invariant variable g.
[…]
In this case, the ergodic decomposition recovers de Finetti’s representation theorem (definetti1929),
$\mathrm{p}(\mathbf{z})=\int_{\mathcal{G}_{{\rm perm}}}\prod_{\omega\in\Omega}\textrm{g}_{\rm perm}(z_{\omega})\,\mathrm{d}\mu(\textrm{g}_{\rm perm}).$ (13)
The invariant variable $\textrm{g}_{\mathrm{perm}}$ is a random probability measure on $\mathcal{Z}$, and $\mu$ is a distribution over such random measures.
[…]
$(z_{ij})\overset{d}{=}(z_{\pi(i),\pi^{\prime}(j)})\quad\text{for all }(\pi,\pi^{\prime})\in\Phi.$ (14)
Under separate exchangeability, the Aldous–Hoover theorem provides the ergodic decomposition (Aldous1981jointy; Aldous1985).
Theorem 2 (Aldous–Hoover representation for separate exchangeability).
A random array $\bm{z}=(z_{ij})_{i,j\in\mathbb{N}}$ is separately exchangeable if and only if its distribution can be written as
$\mathrm{p}(\bm{z})=\int_{\mathcal{G}_{\textrm{sep}}}\mathrm{p}(\bm{z}\,|\,\textrm{g}_{\textrm{sep}})\,\mathrm{d}\mu(\textrm{g}_{\textrm{sep}}).$ (15)
The invariant variable is a random function
$\textrm{g}_{\mathrm{sep}}:[0,1]^{3}\to\mathcal{Z}$ (16)
and the ergodic distribution $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ is defined through uniform random variables
$u_{i},v_{j},u_{ij}$ $\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}(0,1)\quad z_{ij}=g_{\mathrm{sep}}(u_{i},v_{j},u_{ij})\quad\text{for all }i,j\in\mathbb{N}.$ (17)
Through the shared row and column variables $(u_{i})_{i\in\mathbb{N}}$ and $(v_{j})_{j\in\mathbb{N}}$, the entries $z_{ij}$ exhibit dependence within rows and within columns.
Variants of this representation capture other symmetry assumptions. If the array is jointly exchangeable—invariant when rows and columns are permuted together—then the same decomposition holds with
$u_{i},u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(0,1)\quad z_{ij}=\textrm{g}_{\mathrm{joint}}(u_{i},u_{j},u_{ij})\quad\textit{for all }i,j\in\mathbb{N}.$
In a binary array with joint exchangeability, i.e., a random graph, the Aldous–Hoover representation yields
$z_{ij}=\mathbf{1}\{u_{ij}\leq\textrm{g}_{\mathrm{graph}}(u_{i},u_{j})\},$
for a random symmetric graphon $\textrm{g}_{\mathrm{graph}}:[0,1]^{2}\to[0,1]$. In §˜D.3, we discuss array exchangeability conditional on row- and column-specific covariates. Orbanz2015 extends all these representations to $k$-arrays, where $\Omega=\mathbb{N}^{k}$.
Stationary processes. We now consider variables in time or space. Let $\Omega=\mathbb{R}^{d}$ (continuous space) or $\mathbb{Z}^{d}$ (lattice), and $\mathcal{Z}=\mathbb{R}$. A distribution $\mathrm{p}(\bm{z})$ on $\mathcal{Z}^{\Omega}$ is stationary if it is invariant under spatial shifts $\Phi_{\mathrm{shift}}=\{\phi_{h}:\omega\mapsto\omega+h\}$ with $h\in\Omega$:
$\mathrm{p}(\phi_{h}^{-1}(A))=\mathrm{p}(A)\quad\text{for all measurable }A\subseteq\mathcal{Z}^{\Omega},\ h\in\Omega.$
Each such $\mathrm{p}(\bm{z})$ is a stationary probability measure. Among these, a measure $\textrm{g}_{\mathrm{erg}}$ is called ergodic if it cannot be written as a nontrivial mixture of other stationary measures. Equivalently, it is ergodic if $\textrm{g}_{\mathrm{erg}}(A)\in\{0,1\}$ for every shift-invariant event $A$.
By the ergodic decomposition theorem (Shields1996, Theorem I.4.10), any stationary measure can be written uniquely as a mixture of ergodic stationary measures,
$\mathrm{p}(\bm{z})=\int_{\mathcal{G}}\textrm{g}_{\mathrm{erg}}(\bm{z})\,\mathrm{d}\mu(\textrm{g}_{\mathrm{erg}}),$ (18)
where $\mathcal{G}$ is the space of all ergodic stationary measures on $\mathcal{Z}^{\Omega}$. Compared with exchangeable models, stationary processes impose a weaker symmetry: the discrete shift group grows only polynomially with system size, while the permutation group grows super-exponentially (Lubotzky2003). As a result, representation theorems for stationarity are less restrictive, but also less tractable.
2.3 The finite population model
We have defined our model in terms of the entire set $\Omega$. But in practice we only observe finite data. We now specialize our population model to a finite domain.
Let $S_{n}\subseteq\Omega$ be a finite subset, with $S_{1}\subseteq S_{2}\subseteq\cdots$ forming a nested sequence that exhausts $\Omega$. We observe finite data $\bm{x}_{n}=(x_{\omega})_{\omega\in S_{n}}$, each associated with a local latent variable $z_{\omega}$. We write $\bm{z}_{n}=(z_{\omega})_{\omega\in S_{n}}$. Intuitively, this means we observe a finite subset of the full population $\bm{x}$ at selected index locations, i.e. certain entries of an array or certain locations in time or space).
We first state the finite-data analog of the empirical Bayes assumption, that there exists a true prior that captures the population distribution.
Proposition 1 (Finite empirical Bayes model).
Under Assumption˜1, the marginal distribution of the finite data $\bm{x}_{n}$ can be written as
$\mathrm{p}^{\star}(\bm{x}_{n})=\int_{\mathcal{Z}^{S_{n}}}\mathrm{p}^{\star}(\bm{z}_{n})\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\,|\,z_{\omega})\,d\bm{z}_{n},$
for some prior $\mathrm{p}^{\star}(\bm{z}_{n})$ on the latent variables indexed by $S_{n}$.
We next connect this finite prior to the group symmetry, and write its ergodic decomposition.
Proposition 2 (Finite $\Phi$-invariant prior).
Suppose the population prior $\mathrm{p}^{\star}(\bm{z})$ is $\Phi$-invariant as in Assumption˜2. Then its marginal $\mathrm{p}^{\star}(\bm{z}_{n})$ also admits an ergodic representation,
$\mathrm{p}^{\star}(\bm{z}_{n})=\int_{\mathcal{G}}\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})\,d\mu^{\star}(\textrm{g}),$
where
$\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})=\int_{\mathcal{Z}^{\Omega\setminus S_{n}}}\mathrm{p}(\bm{z}\,|\,\textrm{g})\,d\bm{z}_{\Omega\setminus S_{n}}$
is the marginal of the ergodic distribution $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ on $S_{n}$.
Consider a dataset $\bm{x}_{n}$ and assume it was drawn from $\bm{z}_{n}$ with a $\Phi$-symmetric prior. As a consequence of Propositions˜1 and 2, its generative process is a hierarchical model:
g $\sim\mu^{\star}$ (19)
$\bm{z}_{n}$ $\sim\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})$ (20)
$x_{\omega}$ $\sim\mathrm{p}(x\,|\,z_{\omega})\quad\omega\in S_{n}.$ (21)
The latent variables are $\bm{z}_{n}$ and g, and the observations are $\bm{x}_{n}$. The only unknown factor is the prior $\mu^{\star}$, because the likelihood $\mathrm{p}(x\,|\,z)$ is fixed and the form of $\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})$ is known and determined by the probabilistic symmetry.
2.4 Bayesian empirical Bayes
We are given a dataset $\bm{x}_{n}$ and assume a probabilistic symmetry on its prior $\mathrm{p}(\bm{z}_{n})$. Consequently, our data comes from the model in Eq.˜19, with the form of the invariant variable g and ergodic distribution $\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})$ determined by the assumed symmetry. Our goal is to compute the posterior $\mathrm{p}^{\star}(\bm{z}_{n}\,|\,\bm{x}_{n})$.
We first write the posterior as an integral
$\mathrm{p}^{\star}(\bm{z}_{n}\,|\,\bm{x}_{n})=\int_{\mathcal{G}}\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g},\bm{x}_{n})\,\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})\,\mathrm{d}\textrm{g}.$ (22)
Note the conditional distribution $\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g},\bm{x}_{n})$ is proportional to the product of the likelihood $\mathrm{p}(\bm{x}_{n}\,|\,\bm{z}_{n})$ and the ergodic distribution $\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})$:
$\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g},\bm{x}_{n})\propto\mathrm{p}(\bm{x}_{n}\,|\,\bm{z}_{n})\,\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g}),$ (23)
where both these terms are known. The only unknown component in Eq.˜22 is $\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})$, which depends on the unknown $\mu^{\star}$.
We next approximate $\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})$ by a point estimate $\delta_{\hat{\textrm{g}}}$ at its value that maximizes the marginal likelihood of the data,
$\hat{\textrm{g}}\in\operatorname*{arg\,max}_{\textrm{g}\in\mathcal{G}}\log\int_{\mathcal{Z}^{S_{n}}}\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})\,\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\,|\,z_{\omega})\,\mathrm{d}\bm{z}_{n}.$ (24)
The form of $\textrm{g}\in\mathcal{G}$ depends on the chosen symmetry. For example, in exchangeable data, g is a prior on $z$ and $\mathcal{G}$ is the space of such priors. In separately-exchangeable array data, g are the functions described in Eq.˜16; we will parameterize them with neural networks to capture a large space $\mathcal{G}$.
This approximation is motivated by both asymptotic and computational considerations. The observed dataset $\bm{x}_{n}$ is generated from a single draw $\textrm{g}^{\star}$. So asymptotically, we expect the ideal posterior $\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})$ to concentrate at $\textrm{g}^{\star}$ as $n\to\infty$. In particular, by Theorem˜1, the law $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ is $\Phi$-ergodic and hence identifiable (see, e.g., (Kallenberg1997foundations, Lemma 27.1) and (Lindenstrauss2001, Theorem 1.2)). Under a characteristic likelihood, such as the normal model with known variance, distinct priors $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ induce distinct marginals $\mathrm{p}(\bm{x}\,|\,\textrm{g})$, so g is identifiable from $\mathrm{p}(\bm{x}\,|\,\textrm{g})$. Posterior concentration then follows from Doob’s theorem (Ghosal2017, Ch. 6). Computationally, although the marginal likelihood in Eq.˜24 may be intractable, we can use variational inference to approximate it (Blei2017).
Finally, we substitute the point-estimate approximation of the prior into the posterior of Eq.˜22. This gives the approximation
$\hat{\mathrm{p}}(\bm{z}_{n}\,|\,\bm{x}_{n}):=\mathrm{p}(\bm{z}_{n}\,|\,\hat{\textrm{g}},\bm{x}_{n}).$ (25)
[…]
Our procedure is called Bayesian empirical Bayes (BEB). We are given data $\bm{x}_{n}$ and assume a symmetry group $\Phi$ for the latent variables $\bm{z}_{n}$. We use the ergodic representation theorem to form the expanded hierarchical model (Eq.˜19), which includes the invariant variable g and the corresponding ergodic family $\mathrm{p}(\bm{z}\,|\,\textrm{g})$. We parameterize g and estimate it from data via maximum marginal likelihood (Eq.˜24) or a suitable surrogate. Finally, we plug the estimate $\hat{\textrm{g}}$ into the conditional posterior (Eq.˜25). Algorithm˜1 presents this recipe.
Input: Data $\bm{x}_{n}$, likelihood model $\mathrm{p}(\bm{x}_{n}\,|\,\bm{z}_{n})$, group $\Phi$.
0.5em Steps:
  1. 1.
    Derive the ergodic decomposition $\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})$ from $\Phi$.
  2. 2.
    Estimate $\hat{\textrm{g}}$ by maximizing the marginal likelihood or a surrogate:
    $\hat{\textrm{g}}\in\arg\max_{\textrm{g}\in\mathcal{G}}\log\mathrm{p}(\bm{x}_{n};\textrm{g}).$
  3. 3.
    Approximate the posterior of the latent $\bm{z}_{n}$:
[…]
0.5em Output: Approximate posterior $\hat{\mathrm{p}}(\bm{z}_{n}\,|\,\bm{x}_{n})$
Algorithm 1 Bayesian Empirical Bayes
3 New Models for Empirical Bayes
We now use BEB to design new methods for empirical Bayes inference. In our first example, we use BEB to recover the classical empirical Bayes g-modeling for the normal means model (§˜3.1). This recovers classical EB methods (Efron2015deconvolution; Soloff2021). We then extend EB to three additional probabilistic symmetries: separate/joint exchangeability in arrays (§˜3.2), stationarity in spatial data (§˜3.3), and relative exchangeability conditional on covariates (§˜D.3). In each case, we follow the recipe of Algorithm˜1. We start with a dataset, a symmetry group, and a likelihood. We then design an empirical Bayes model–tailored to the symmetry–and approximate the posterior of the latent structured variables.
3.1 Classical Example Revisited: Normal Means Model
Let $\Omega=\mathbb{N}$ and $\mathcal{X}=\mathbb{R}^{d}$. The observed variables $\bm{x}\in\mathcal{X}^{\Omega}$ are generated from the heteroskedastic normal means model
$\bm{z}\sim\mathrm{p}^{\star}(\bm{z}),\quad x_{i}\sim\mathcal{N}\left(z_{i},\Sigma_{i}\right),$ (26)
where $\Sigma_{i}\in\mathbb{R}^{d\times d}$ is known.
We assume the true prior $\mathrm{p}^{\star}(\bm{z})$ is exchangeable. By de Finetti’s theorem, we can write Eq.˜26 as the following generative process:
$\textrm{g}_{\rm perm}\sim\mu^{\star},\quad z_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{g}_{\rm perm},\quad x_{i}\sim\mathcal{N}\left(z_{i},\Sigma_{i}\right).$ (27)
Here $g_{\rm perm}$ is a random prior over $z$. In short, de Finetti’s representation specializes Theorem˜1 to sequences, with $\prod_{i=1}^{\infty}\textrm{g}_{\rm\textrm{perm}}(z_{i})$ the ergodic distribution.
Under the expanded model, the posterior $\mathrm{p}^{\star}(\bm{z}_{n}\mid\bm{x}_{n})$ has the form
$\mathrm{p}^{\star}(\bm{z}_{n}\mid\bm{x}_{n})=\int_{\mathcal{G}}\prod_{i=1}^{n}\mathrm{p}(z_{i}\mid\textrm{g}_{\rm perm},x_{i})\,\mathrm{p}^{\star}(\textrm{g}_{\rm perm}\mid\bm{x}_{n})\,\mathrm{d}\textrm{g}_{\rm perm}.$ (28)
In Bayesian EB, we approximate $\mathrm{p}^{\star}(\textrm{g}_{\rm perm}\mid\bm{x}_{n})$ with its MMLE,
$\hat{\textrm{g}}_{\rm perm}\in\operatorname*{arg\,max}_{\textrm{g}_{\rm perm}\in\mathcal{P}(\mathbb{R}^{d})}\log\int_{\mathcal{X}^{n}}\prod_{i=1}^{n}\left(\mathrm{p}(x_{i}\mid z_{i})\textrm{g}_{\rm perm}(z_{i})\right)\mathrm{d}\bm{z}_{n}.$ (29)
If we restrict $g_{\rm perm}$ to a parametric family, then this recovers parametric empirical Bayes estimation, or type-II maximum likelihood. If we consider all possible priors, then it recovers the nonparametric maximum likelihood estimation (NPMLE) approach.
With $\hat{g}_{\rm perm}$ in hand, we approximate the target posterior from Eq.˜28,
$\mathrm{p}^{\star}(\bm{z}_{n}\mid\bm{x}_{n})\approx\prod_{i=1}^{n}\mathrm{p}(z_{i}\mid x_{i},\hat{\textrm{g}}_{\rm perm}),$
where each posterior is
$\mathrm{p}(z_{i}\,|\,x_{i},\hat{\textrm{g}}_{\rm perm})\propto\mathrm{p}(x_{i}\mid z_{i})\hat{\textrm{g}}_{\rm perm}(z_{i}).$
NPMLE and the normal means model have been thoroughly studied in numerous works (Kiefer1956; James1960; Jiang2009; Brown2009; Jiang2010; Xie2012EB; Efron2012; Efron2015deconvolution; Tan2016; Brown2018; Jiang2020; chen2022empirical; Soloff2021).
If the likelihood is non-Gaussian, the same reasoning yields other nonparametric EB g-modeling procedures. For example, with exponential data $x_{i}\mid z_{i}\sim\mathrm{Exp}(z_{i})$, this approach recovers g-modeling for the exponential sequence model (Robbins1980exponential). As another example, when $x_{i}$ is a sample variance and $z_{i}$ the true variance, we obtain the $\chi^{2}$ sequence model $x_{i}\mid z_{i}\sim(z_{i}/k)\,\chi^{2}_{k}$ (for known $k$). The exchangeability assumption on $z_{1},\cdots,z_{n}$ recovers the g-modeling setting studied in ignatiadis2025empirical.
Denoising an MNIST digit.
We now introduce a running example to illustrate the performance of Bayesian EB methods. The dataset is an image of "5" from the MNIST dataset of handwritten digits (lecun2002gradient). The $28\times 28$ image is corrupted by i.i.d. Gaussian noise with variance equal to the empirical pixel variance of the clean image
$x_{\omega}\sim\mathcal{N}(z_{\omega}^{\star},\sigma_{0}^{2}),\qquad\omega\in[28]^{2}.$
The goal is to recover the clean image $\bm{z}^{\star}_{n,n}$ from the observed noisy image $\bm{x}_{n,n}$. The maximum likelihood estimator is the simply the noisy image which has a MSE of $10.291$. The exchangeable sequence approach treats the problem as a simple normal means problem with known variance $\sigma_{0}^{2}$ and applies the NPMLE procedure. The resulting MSE is 3.613.61–a substantial reduction from the error of the observed image.
3.2 Bayesian empirical Bayes for matrix recovery
Let $\Omega=\mathbb{N}^{2}$ and $\mathcal{X}=\mathbb{R}$. We observe a noisy array $\bm{x}_{n,p}=(x_{ij})_{i\in[n],\,j\in[p]}$ with latent signal $\bm{z}_{n,p}=(z_{ij})_{i\in[n],\,j\in[p]}$ generated by
$x_{ij}\sim\mathcal{N}(z_{ij},\,\tau_{ij}^{-1}),\qquad\tau_{ij}>0.$
Our goal is to estimate $\bm{z}_{n,p}$ from $\bm{x}_{n,p}$. We will use BEB to fit a structured prior that captures dependencies among entries.
We assume $\bm{z}$ is separately exchangeable: its distribution is invariant to independent permutations of rows and columns. Recall by the Aldous–Hoover theorem (Theorem˜2), we can write the conditional distribution $\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g}_{\textrm{sep}})$ as
$u_{i},v_{j},u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\qquad z_{ij}=\textrm{g}_{\textrm{sep}}(u_{i},v_{j},u_{ij}),$ (30)
where the invariant variable is a function $\textrm{g}_{\textrm{sep}}:[0,1]^{3}\to\mathbb{R}$. This conditional representation fully characterizes the finite-dimensional marginals of the ergodic distribution $\mathrm{p}(\bm{z}\mid\textrm{g}_{\textrm{sep}})$, in the context of Theorem˜1.
The corresponding Bayesian empirical Bayes model is
$\textrm{g}_{\textrm{sep}}$ $\sim\mu^{\star},\quad$ (31)
$u_{i},v_{j},u_{ij}$ $\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\quad\quad\quad i\in[n],j\in[n]$ (32)
$z_{ij}$ $=\textrm{g}_{\textrm{sep}}(u_{i},v_{j},u_{ij}),\quad$ (33)
$x_{ij}$ $\sim\mathcal{N}(z_{ij},\,\tau_{ij}^{-1}).$ (34)
We parameterize $\textrm{g}_{\textrm{sep}}$ by a neural network $\textrm{g}_{\textrm{sep},\theta}\in\mathcal{G}_{\mathrm{NN}}$, a two–layer ReLU feed-forward network mapping $0,1]^{3}$ to $\mathbb{R}$. We estimate $\theta$ by marginal maximum likelihood, now writing the conditional distribution of $x_{ij}$ directly in terms of $\bm{u}_{n},\bm{v}_{p},\bm{u}_{n,p}$,
$\hat{\theta}\in\operatorname*{arg\,max}_{\theta\in\Theta}\,\log\int\int\int\prod_{i=1}^{n}\prod_{j=1}^{p}\mathrm{p}\!\left(x_{ij}\mid\textrm{g}_{\textrm{sep},\theta}(u_{i},v_{j},u_{ij})\right)\mathrm{d}\bm{u}_{n}\mathrm{d}\bm{v}_{p}\mathrm{d}\bm{u}_{n,p}.$ (35)
Exact evaluation of Eq.˜35 is intractable because the integral is high–dimensional. We approximate it by maximizing an evidence lower bound (ELBO) under a structured variational family. The estimate $\hat{\textrm{g}}_{\textrm{sep},\theta}$ is the variational MMLE. See Appendix˜D for details.
Given $\hat{\textrm{g}}_{\textrm{sep},\theta}$, we approximate the empirical Bayes posterior for $\bm{z}_{n,p}$ with samples, through the posterior of $\bm{u}_{n},\bm{v}_{p},\bm{u}_{n,p}$. For each sample $s$,
$\bm{u}_{n}^{(s)},\bm{v}_{p}^{(s)},\bm{u}_{n,p}^{(s)}$ $\sim\mathrm{p}(\bm{u}_{n},\bm{v}_{p},\bm{u}_{n,p}\,|\,\bm{x}_{n,p},\hat{\textrm{g}}_{\textrm{sep}})$
$z_{ij}^{(s)}$ $=\hat{\textrm{g}}_{\textrm{sep}}(u_{i}^{(s)},v_{j}^{(s)},u_{ij}^{(s)}).$
The conditional posterior of the auxiliary latent variables is
$\mathrm{p}(\bm{u}_{n},\bm{v}_{p},\bm{u}_{n,p}\,|\,\bm{x}_{n,p},\hat{\textrm{g}}_{\textrm{sep}})\propto\prod_{i=1}^{n}\prod_{j=1}^{p}\mathcal{N}\!\left(x_{ij};\hat{\textrm{g}}_{\textrm{sep}}(u_{i},v_{j},u_{ij}),\tau_{ij}^{-1}\right),$ (36)
which we sample from using the No–U–Turn Sampler (NUTS; Hoffman2014) or a variational approximation.
Related symmetries.
If the array is jointly exchangeable then it is invariant when rows and columns are permuted together. In this case, Eq.˜30 simplifies to
$u_{i},u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\qquad z_{ij}=\textrm{g}_{\textrm{joint}}(u_{i},u_{j},u_{ij}),$
with $\textrm{g}_{\textrm{joint}}:[0,1]^{3}\to\mathbb{R}$. We follow nearly the same procedure.
If the array is relatively exchangeable conditioned on row covariates $\bm{y}_{n}=(y_{i})_{i=1}^{n}$ with $y_{i}\in\mathcal{Y}$ and column covariates $\bm{a}_{p}=(a_{j})_{j=1}^{p}$ with $a_{j}\in\mathcal{A}$, then Eq.˜30 expands to
$u_{i},v_{j},u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\qquad z_{ij}=\textrm{g}_{\textrm{rel}}\bigl(y_{i},a_{j},u_{i},v_{j},u_{ij}\bigr),$
with $\textrm{g}_{\textrm{rel}}:\mathcal{Y}\times\mathcal{A}\times[0,1]^{3}\to\mathbb{R}$. This yields the covariate-assisted Bayesian empirical Bayes procedure developed in §˜D.3.
If the array is fully exchangeable then it is invariant to shuffling all of its cells. This case reduces to the algorithms of §˜3.1, where $z_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{g}_{\textrm{perm}}$ for a random measure $\textrm{g}_{\textrm{perm}}$.
Denoising an MNIST digit (continued).
We revisit the example of denoising an MNIST "5" from Figure˜1. The image is treated as a noisy matrix and we apply the empirical Bayes matrix recovery under separate and joint exchangeability. We report $100\times\textrm{MSE}$ between the posterior means $\hat{\bm{z}}_{n,n}$ and the true pixel value $\bm{z}^{\star}_{n,n}$. The fourth and fifth panels in Figure˜1 show the results. EBMR with a jointly exchangeable prior (EBMR–Joint) achieves the lowest error, followed by EBMR with a separately exchangeable prior (EBMR–Sep), then EBMR with a fully exchangeable prior (NPMLE). This ordering aligns with the symmetry hierarchy full $\Rightarrow$ separate $\Rightarrow$ joint: weaker (less restrictive) symmetry assumptions enlarge the class of admissible priors and increase flexibility. On the other hand, weaker symmetry assumptions also increase the sample and computational complexity for estimating g. This creates a trade-off. In our setting, joint exchangeability offers a sweet spot between expressiveness and efficiency.
3.3 Bayesian empirical Bayes for spatial regression
Let $\Omega=\mathbb{R}^{d}$ and $\mathcal{X}=\mathbb{R}$. We observe $\bm{x}_{n}=\{x_{\omega}:\omega\in S_{n}\}$ at sites $S_{n}\subset\Omega$. Each site has covariates $a_{\omega}=(a_{\omega,1},\ldots,a_{\omega,p})\in\mathbb{R}^{p}$. The spatial regression model is
$x_{\omega}=\beta^{\top}a_{\omega}+z_{\omega}+\epsilon_{\omega},\qquad\epsilon_{\omega}\sim\mathcal{N}(0,\tau_{\omega}^{-1}),\ \tau_{\omega}>0,$ (37)
where $\beta\in\mathbb{R}^{p}$ are regression coefficients and $\bm{z}=(z_{\omega})_{\omega\in\Omega}$ is a latent process that captures spatial dependence. We place a flat prior $\pi(\beta)\propto 1$, so the estimator of $\beta$ coincides with weighted least squares. The resulting model along with Eq.˜37 coincide with the standard model in Bayesian spatial regression (handcock1993bayesian; Stein1999; DeOliveira1997spatial; Berger2001spatial; Banerjee2008; Banerjee2014spatial; Li2024spatial).
We assume $\bm{z}$ is stationary: its distribution is invariant under translations $\phi_{h}:\omega\mapsto\omega+h$. Following the ergodic decomposition in Eq.˜18, any stationary process can be written as a mixture of ergodic stationary measures,
$g\sim\mu^{\star},\qquad\bm{z}\mid g\sim g,$ (38)
where $g$ is a stationary probability measure on $\mathbb{R}^{\Omega}$. Following the BEB method, our goal is to estimate the invariant variable $g$ from data.
Defining stationary measures with Gaussian processes.
The space of stationary measures is complex and so, in practice, we restrict the ergodic family $\mathcal{G}$ to stationary ergodic Gaussian processes. A Gaussian process $\mathrm{GP}(0,k)$ is stationary if its covariance function depends only on the displacement $k(\omega,\omega^{\prime})=f(\omega-\omega^{\prime}):=k(\Delta)$. Each stationary covariance function $k$ with an atomless spectral density (as described below) defines an ergodic stationary Gaussian measure on $\mathbb{R}^{\Omega}$ (maruyama1949harmonic). So we can write
$k\sim\mu(k),\qquad\bm{z}\mid k\sim\mathrm{GP}(0,k),$ (39)
where $\mu(k)$ is a distribution over stationary kernels.
How can we characterize the stationary covariance functions? Bochner’s theorem (bochner2005harmonic) states that $k$ is stationary and positive definite if and only if it can be written as
$k(\Delta)=\int_{\mathbb{R}^{d}}e^{\,2\pi i\,s^{\top}\Delta}\,\psi(s)\,ds,$ (40)
for some nonnegative integrable function $\psi$. The function $\psi:\mathbb{R}^{d}\to[0,\infty)$ is called the spectral density. It describes how the variance of the process is distributed across spatial frequencies. The associated spectral measure is atomless if it assigns zero mass to every singleton, i.e., $\Psi({s})=0$ for all $s\in\mathbb{R}^{d}$ (in particular, if $\Psi$ has density $\psi$, it is atomless). Bochner’s theorem gives a one–to–one correspondence between stationary Gaussian measures and their spectral densities. We therefore identify each ergodic stationary measure $g$ with its spectral density $\psi$:
$\psi\sim\mu^{\star},\qquad\bm{z}\mid\psi\sim\mathrm{GP}(0,k_{\psi}),$ (41)
where $k_{\psi}$ is the kernel defined by (40). The empirical Bayes task is to estimate $\psi$ from the marginal distribution of the observed data.
Bayesian empirical Bayes for spatial regression.
We have defined the invariant variable $\bm{g}$ as the spectral density of a stationary kernel for a Gaussian process. With this definition, the BEB model is
$\bm{z}$ $\sim\mathrm{GP}(0,k_{\psi}),$ $\beta$ $\sim\pi(\beta)\propto 1,$ $x_{\omega}$ $\sim\mathcal{N}(\beta^{\top}a_{\omega}+z_{\omega},\tau_{\omega}^{-1}).$ (42)
We estimate the invariant variable $\psi$ by maximizing the marginal likelihood,
$\hat{\psi}\in\operatorname*{arg\,max}_{\psi\in\mathcal{G}}\log\int\Bigl[\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\mid z_{\omega},\beta)\Bigr]\mathrm{p}(\bm{z}_{n}\mid\psi)\pi(\beta)\,d\beta\,d\bm{z}_{n}.$ (43)
[…]
In practice, we model $\psi$ as a mixture of Gaussians, and fit it by maximum marginal likelihood. The details of this procedure, as well as how we perform simultaneous posterior inference, are given in Appendix˜E.
In the spatial regression example, we model the spectral density of the Gaussian-process kernel as a mixture of Gaussian distributions. Under this model, the empirical Bayes procedure reduces to selecting the GP kernel parameters by maximum likelihood (Rasmussen2006). This approach is a common practice in the GP literature. It comes with strong theoretical guarantees and retains the conjugacy of the GP posterior (Loh2000; Bachoc2014; Szab2015; Xu2017). What we see here is that this approach can be interpreted as an instance of Bayesian empirical Bayes.
Denoising an MNIST digit (continued).
The image of a handwritten digit does not change its meaning if we shift the image, so stationarity is a reasonable assumption. We fit EB spatial regression with a constant covariate ($a_{\omega}\equiv 1$) and a $K=10$ Gaussian–mixture spectral density, using the inference and sampling steps outlined in Appendix˜E. The rightmost panel of Figure˜1 shows that EB spatial regression denoises better than NPMLE but not as well as EBMR–sep or EBMR–joint. Nonetheless, all EB methods substantially improve over the MLE. Moreover, all BEB methods that use symmetries beyond exchangeability improve over the NPMLE.
Figure 1: EB spatial regression on a noisy MNIST “5.” Top: original (left), noisy (middle), and NPMLE (right). Bottom: EBMR–sep (left), EBMR–joint (middle), and EB spatial regression (right). Errors are $100\times\textrm{MSE}$ of posterior means; EBMR with a jointly exchangeable prior performs best, followed by EBMR with a separately exchangeable prior and EB spatial regression, then NPMLE.
4 Simulation Studies
We evaluate the new Bayesian empirical Bayes methods using synthetic data.
  • In §˜4.1, we assess empirical Bayes matrix recovery (EBMR; §˜3.2) across 60 settings by varying sample size, noise level, and the data-generating mechanism; we implement EBMR with a separately exchangeable prior (EBMR–Sep) and compare against NPMLE and empirical Bayes matrix factorization (EBMF; Wang2021EBMF).
  • In §˜4.2, we evaluate the covariate-assisted empirical Bayes (CAEB; §˜D.3) method for denoising matrices with row- and column-specific covariates and compare it to EBMR–Sep, EBMF, and NPMLE.
  • In §˜4.3, we apply empirical Bayes spatial regression (spatial EB; §˜3.3) to synthetic time-series data. We show that the MSE of EB spatial regression vanishes with increasing sample size, whereas the NPMLE plateaus at a constant risk.
In each study, we generate observations $\bm{x}_{S_{n}}$ on a finite domain $S_{n}$ (e.g. $S_{n}=[n]$) from a known data-generating process with latent truth $\bm{z}^{\star}_{S_{n}}$. We report the relative MSE to assess how well the BEB posterior mean $\hat{\bm{z}}_{S_{n}}$ recovers the true latent values $\bm{z}^{\star}_{S_{n}}$. The relative MSE measures the MSE of the estimator as a percentage of the MSE of the MLE (out of $100$):
$\textrm{R-MSE}=\frac{1}{\left|S_{n}\right|}\sum_{\omega\in S_{n}}\left(\hat{z}_{\omega}-z_{\omega}^{\star}\right)^{2}\times 100\times\tau,$
where $\tau$ is the precision of the noise distribution.
A better method achieves lower R-MSE.
4.1 Empirical Bayes Matrix Recovery
In our first study, we simulate observations from the normal means model $x_{ij}\sim\mathcal{N}\left(z_{ij}^{\star},\tau^{-1}\right)$ with known precision $\tau>0$. The true matrix $\bm{z}_{n,p}^{\star}$ is a single draw from the Aldous–Hoover representation with a pre-specified function $T_{0}$, that is
$u_{i},v_{j},u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\qquad z_{ij}^{\star}=T_{0}\left(u_{i},v_{j},u_{ij}\right).$
We consider five choices of $T_{0}$, and vary the precisions $\tau\in\{0.25,1,2,4\}$ and the sample sizes $(n,p)\in\{(20,20),\ (20,50),\ (50,50)\}$. The different $T_{0}$ are as follows:
  • Linear: $T_{0}(U,V,W)=U+V+W$;
  • Sine–log: $T_{0}(U,V,W)=\sin(\pi UV)+\log(1+W)$;
  • Sine–cos: $T_{0}(U,V,W)=\dfrac{\sin(\pi U)\cos(\pi V)}{1+W^{2}}$;
  • Tanh: $T_{0}(U,V,W)=\tanh(U+V+W)$;
  • Reciprocal: $T_{0}(U,V,W)=\dfrac{1}{1+\lvert U+V+W\rvert}$.
We examine all combinations of $(\tau,n,p,T_{0})$ for a total of $60$ settings. For each setting, we run the experiment 10 times to calculate the R-MSE. We compare EBMR with a separately exchangeable prior (EBMR–Sep, Algorithm˜2) against three benchmarks:
  • Benchmark 1: MLE. The MLE is $\hat{z}_{\omega}=x_{\omega}$; by construction its R-MSE equals $100$.
  • Benchmark 2: NPMLE. NPMLE-based methods are state of the art for normal-means denoising problems (Jiang2009; Koenker2014; Soloff2021). In the context of matrix recovery, it is equivalent to EBMR with a fully exchangeable prior, $z_{\omega}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{g}$. We estimate g via the nonparametric MLE using sequential quadratic programming (Kim2020NPMLE), implemented with ebnm in R.
  • Benchmark 3: Empirical Bayes Matrix Factorization (EBMF). EBMF (Wang2021EBMF) models the latent matrix as low rank with exchangeable loadings and factors,
    $\bm{z}_{n,p}=\sum_{k=1}^{C}\ell_{k}f_{k}^{\top},\qquad\ell_{k}\in\mathbb{R}^{n},\ f_{k}\in\mathbb{R}^{p},$
    with $\{\ell_{k}\}_{k=1}^{C}\stackrel{{\scriptstyle iid}}{{\sim}}$ from an unknown prior $\textrm{g}_{\ell}$ and $\{f_{k}\}_{k=1}^{C}\stackrel{{\scriptstyle iid}}{{\sim}}$ from an unknown prior $\textrm{g}_{f}$. The priors $(\textrm{g}_{\ell},\textrm{g}_{f})$ are fit by variational empirical Bayes under the normal-means likelihood for $\bm{x}_{n,p}$, and we use the posterior mean of $\bm{z}_{n,p}$ as the estimator. Unless noted otherwise, we set $C=10$ and implement the method using the flashier package in R.
To implement Algorithm˜2, we set $K=10$ and parameterize $\textrm{g}_{\textrm{sep}}:[0,1]^{3}\mapsto\mathbb{R}$ as a fully connected feed-forward MLP with two hidden layers, ReLU activations, and 5 units per layer. We choose a simple network configuration because the sample sizes are small.
Table 1 shows the results that EBMR–Sep outperforms NPMLE and EBMF in $55$ out of the $60$ settings. When $T_{0}$ introduces nonlinear dependencies across rows and columns (e.g., the sine–log function), the error reduction of EBMR–Sep relative to NPMLE and EBMF is substantial. The few cases where EBMF outperforms EBMR–Sep appear for small sample sizes—e.g., (n,p)=(20,20)(n,p)=(20,20)—in settings where the latent factors $(u_{i},v_{j})$ interact via an approximate low-rank structure (as with the sine–cos function).
Table 1: EBMR-Sep attains the lowest R-MSE in 55 of 60 settings. Matrix-recovery performance for NPMLE, EBMR with a separately exchangeable prior (Sep), and EBMF across simulated settings. Reported values are the median relative MSE (R-MSE), defined as each method’s MSE as a percentage of the MSE of the MLE. The lowest R-MSE in each setting is shown in bold.
n p $\bm{\tau}$ Linear Sine–log Sine–cos Tanh Reciprocal NPMLE Sep EBMF NPMLE Sep EBMF NPMLE Sep EBMF NPMLE Sep EBMF NPMLE Sep EBMF 20 20 0.1 3.23 3.31 9.16 2.09 1.90 9.24 2.85 2.33 2.14 1.08 0.60 7.57 0.88 0.53 1.84 20 20 0.25 7.28 6.45 10.15 3.83 3.50 9.95 5.64 4.83 5.36 0.96 0.58 8.63 0.75 0.43 4.61 20 20 1 23.35 15.99 18.56 13.47 11.93 14.83 18.74 10.01 9.24 2.87 2.22 9.92 1.83 1.56 9.91 20 20 4 53.24 34.80 43.95 36.01 26.62 31.50 48.74 14.57 13.19 8.93 7.16 13.07 4.65 4.27 10.72 20 50 0.1 2.88 2.76 8.26 2.00 1.73 10.38 2.49 2.27 2.06 0.65 0.39 7.75 0.61 0.49 1.75 20 50 0.25 6.38 5.11 9.79 4.17 3.55 8.27 5.31 3.54 5.14 0.87 0.47 5.68 0.70 0.30 4.38 20 50 1 20.13 12.47 22.78 14.11 9.04 18.03 17.94 6.22 20.55 2.29 1.42 8.05 1.73 0.92 7.35 20 50 4 49.97 29.56 69.45 38.82 22.01 49.92 44.95 8.67 37.59 5.55 4.77 10.64 3.15 3.00 9.00 50 50 0.1 2.69 2.24 4.73 1.69 1.45 4.00 1.88 1.22 1.84 0.37 0.27 3.65 0.25 0.20 1.95 50 50 0.25 6.21 4.30 6.00 3.86 3.06 5.13 4.49 2.12 4.61 0.65 0.61 4.05 0.38 0.32 4.83 50 50 1 20.65 10.76 12.23 13.54 7.49 8.63 15.73 2.76 4.48 2.27 1.98 4.84 1.24 1.01 4.32 50 50 4 50.99 28.45 37.36 38.32 19.66 23.11 42.34 4.85 6.28 8.07 5.77 7.77 4.03 3.32 5.09
4.2 Covariate-Assisted Empirical Bayes
We evaluate the posterior-mean estimator of covariate-assisted empirical Bayes (CAEB; §˜D.3) for matrix recovery with side information. Data are generated according to
$z_{ij}^{\star}=T_{0}(y_{i},a_{j},u_{i},v_{j},w_{ij}),\qquad x_{ij}\sim\mathcal{N}(z_{ij}^{\star},1),$
where $y_{i}\in\mathbb{R}^{3}$ and $a_{j}\in\mathbb{R}^{4}$ are global row- and column-specific covariates, $u_{i}$ and $v_{j}$ are latent factors, and $w_{ij}$ is an additional stochastic input.
We simulate $y_{i}$ as iid draws from a standard $t$-distribution with $5$ degrees of freedom and $a_{j}$ from a beta distribution with $a=2$ and $b=5$. We consider three choices of $T_{0}$:
  • Linear: $T_{0}(y,a,u,v,w)=\sum_{k=1}^{3}y[k]+\sum_{k=1}^{4}a[k]+uv+w$
  • Nonlinear: $T_{0}(y,a,u,v,w)=\sum_{k=1}^{3}y[k]+\sum_{k=1}^{4}a[k]+\sin(\pi u)\cos(\pi v)+0.5w^{2}$
  • Logistic: $T_{0}(y,a,u,v,w)=\sum_{k=1}^{3}y[k]+\sum_{k=1}^{4}a[k]+\text{sigmoid}(u+v)+w$
where $y[k]$ and $a[k]$ denote the $k$th coordinates of $y\in\mathbb{R}^{3}$ and $a\in\mathbb{R}^{4}$.
Implementation of covariate-assisted EB. We compare CAEB, EBMR–Sep, NPMLE, EBMF, and MLE. To implement CAEB, we parameterize $\textrm{g}_{\textrm{rel}}:\mathbb{R}^{7}\times[0,1]^{3}\mapsto\mathbb{R}$ as a two-layer ReLU neural network with 10 hidden units per layer. We perform the stochastic gradient step using Adam and use NUTS with a target acceptance probability of $0.8$ to sample from the posterior.
Table 2 reports the median R-MSE of the competing methods. CAEB consistently outperforms the NPMLE. In some cases, covariate-assisted EB achieves $10\%$$20\%$ of the R-MSE of the NPMLE method. Thus, for simulated matrix data with covariates, covariate-assisted EB provides a markedly better denoising procedure than NPMLE.
Table˜2 also shows that Bayesian EB methods achieve lower R-MSE than NPMLE. For smaller matrices $(n,p)\in\{(20,20),(20,50)\}$, EBMR–Sep beats CAEB. The reason is that estimating the 10-dimensional map $\textrm{g}_{\textrm{rel}}$ requires more data than the 3-dimensional $\textrm{g}_{\textrm{sep}}$. Once $n,p\geq 50$, CAEB overtakes EBMR–Sep. At $n=p=100$, the sample size suffices to fit the 10-dimensional network for $\textrm{g}_{\textrm{rel}}$, and the benefit of covariates becomes evident. Figure˜2 further shows that with $n=100$ and $p$ increasing from $100$ to $1000$, CAEB dominates across all three generative settings.
$n$ $p$ Linear Logistic Nonlinear
CAEB SEP NPMLE EBMF CAEB SEP NPMLE EBMF CAEB SEP NPMLE EBMF
20 20 50.63 18.79 77.19 24.66 44.45 18.63 82.69 20.83 42.76 17.55 81.10 20.88
20 50 29.25 15.69 75.59 28.82 27.23 13.93 75.71 27.20 28.37 14.74 79.83 43.26
20 100 20.21 12.92 76.71 23.27 17.95 12.79 76.27 22.40 14.71 12.15 78.65 31.35
50 50 19.19 13.02 80.47 16.50 17.65 12.62 80.19 15.30 13.86 11.34 81.24 10.82
50 100 14.66 13.02 82.64 17.87 12.90 11.32 77.83 15.69 9.94 10.42 79.03 11.71
100 100 10.18 11.26 81.93 12.26 9.04 10.52 80.97 11.84 5.81 10.24 83.67 7.52
[…]
Figure 2: Covariate-assisted EB achieves the lowest risk in large-sample settings. When $p$ is large, all methods achieve constant MSE that reflects the oracle risk as determined by the symmetry group. Covariate-assisted EB achieves the lowest risk compared to other benchmarks.
We also assess the robustness of BEB methods to misspecification of the noise process. Specifically, we rerun the simulations with noise drawn from a $t$-distribution with degrees of freedom $5$ and scale parameter $1$ instead of the $N(0,1)$ distribution. While the absolute MSEs of the methods are slightly higher, the qualitative comparison among the methods is unchanged. See Figure˜8 in the Appendix for details.
4.3 Empirical Bayes Spatial Regression
Finally, we evaluate empirical Bayes spatial regression with simulated data. We generate spatial data on an evenly spaced grid $10=\omega_{1}<\omega_{2}<\cdots<\omega_{n}=10$. The observed variables follow the model
$x_{\omega_{i}}=a_{\omega_{i}}^{\top}\beta^{\star}+z_{\omega_{i}}^{\star}+\varepsilon_{\omega_{i}},\qquad i=1,\ldots,n,$
with design vectors $a_{\omega_{i}}=(1,\xi_{\omega_{i}})\in\mathbb{R}^{p}$ where $\xi_{\omega_{i}}\sim\mathcal{N}(0,I_{p-1})$, and noise $\varepsilon_{\omega_{i}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau^{-1})$.
We fix $p=3$ and set $\beta^{\star}=(0.5,-1.2,0.3)^{\top}$. The latent field $z^{\star}=(z_{\omega_{1}}^{\star},\ldots,z_{\omega_{n}}^{\star})$ is a stationary Gaussian process with covariance kernel
$\Sigma^{\star}_{ij}=k_{w,\sigma,\mu}\left(\omega_{i}-\omega_{j}\right),\quad w=(0.5,0.3,0.2),\quad\sigma=(1.0,0.2,0.35),\quad\mu=(0,0.15,-0.25),$
where the formula for $k_{w,\sigma,\mu}$ is in Eq.˜69. Equivalently, it is a stationary kernel with a three-component Gaussian mixture $0.5\mathcal{N}(0,1^{2})+0.3\mathcal{N}(0.15,0.2^{2})+0.2\mathcal{N}(-0.25,0.35^{2})$ as the spectral density. We draw a single $\bm{z}_{n}^{\star}\sim\mathcal{N}(0,\Sigma^{\star})$ as the ground truth.
To simulate $\bm{x}_{n}$, we vary the precision $\tau\in\{0.25,1,4\}$ and the sample size $n$ from $100$ to $3000$. We fit spatial EB regression (§˜3.3) with $K\in\{3,4,5\}$, draw posterior samples of $\beta$ and $\bm{z}_{n}$, and report the posterior mean of $\bm{z}_{n}$ and its R-MSE.
As a baseline, we apply the NPMLE to residuals under the true coefficient $\beta^{\star}$:
$r_{\omega}=x_{\omega}-a_{\omega}^{\top}\beta^{\star},\qquad r_{\omega}\sim\mathcal{N}\bigl(z_{\omega}^{\star},\ \tau_{\omega}^{-1}\bigr),$
and estimate $\bm{z}_{n}^{\star}$ using the normal-means model with a fully exchangeable prior (§˜3.1). The MLE for $z_{\omega}$ in this case is $r_{\omega}$.
Figure˜3 shows the results. EB spatial regression uniformly dominates the regression-adjusted NPMLE, which indicates that stationarity is a more informative symmetry than full exchangeability for this example. At each noise level, the MSE of the EB spatial estimator of $\bm{z}_{n}^{\star}$ decreases sharply with the sample size, while the MSE of the regression-adjusted NPMLE remains essentially flat as $n$ increases. The performance of the two methods is comparable only for small samples (e.g., $n=100$); for larger $n$, the spatial EB estimator achieves substantially lower error. In Figure˜10 of Appendix˜G, we show that the MSE of the posterior-mean estimator for $\beta$ also decreases with $n$ under the spatial EB model, which suggests that the posterior of $\beta$ satisfies posterior consistency.
In high-precision regimes, the NPMLE yields little improvement, with its R-MSE hovering around $80$. By contrast, the R-MSE of EB spatial regression drops below $20$ once $n\geq 2000$, albeit with higher cross-replicate variability than in lower-precision regimes.
Figure 3: EB spatial regression dominates NPMLE. Relative MSE for recovering the latent series $\bm{z}_{n}^{\star}$ across sample size $n$ and noise precision $\tau$. EB spatial regression is fit with $K\in\{3,4,5\}$ mixtures and is stable to modest over-specification of $K$. The NPMLE baseline uses residuals $r_{\omega}=x_{\omega}-a_{\omega}^{\top}\beta^{\star}$ with $r_{\omega}\sim\mathcal{N}(z_{\omega},\tau_{\omega}^{-1})$ and a fully exchangeable prior. Points show medians over 10 replicates; bands show the interquartile range (25–75%). EB error decreases with $n$, whereas NPMLE error remains roughly flat.
5 Real Data Examples
For applications, we consider (i) a semisynthetic study where we add Gaussian noise to a cancer gene–expression matrix with varying noise levels, and (ii) denoising and interpolating NYC air–quality data using EB spatial regression. In the first example, we evaluate our methods using R–MSE of the recovered matrix versus the ground truth. In the second example, we provide the full inference results to reason about the learned prior and posterior.
5.1 Cancer
We apply empirical Bayes matrix recovery methods to a gene–expression dataset under varying levels of noise corruption. The dataset is a $251\times 226$ gene–expression matrix analyzed by Carvalho2008 and Knowles2011, containing measurements of 226 genes across 251 breast–cancer patients. The statistical goal is to recover the latent gene–expression matrix from a noisy observation. It is reasonable to assume a priori that genes are exchangeable and patients are exchangeable, so we apply the empirical Bayes matrix recovery method with a separately exchangeable prior to pool information across related gene-specific and patient-specific effects.
In the semisynthetic design, we add independent Gaussian noise with precision $\tau\in\{0.01,0.03,0.1,0.3,1,3,10\}$ to each entry and compare three empirical Bayes procedures: (i) EBMR with a separately exchangeable prior (EBMR–sep), (ii) nonparametric maximum likelihood estimation (NPMLE; Jiang2009), and (iii) empirical Bayes matrix factorization (EBMF; Wang2021EBMF). We measure the performance by the relative mean–squared error (R–MSE) between the recovered and original matrices, averaged over 20 replicates and summarized by the median R–MSE.
For EBMF, we follow Wang2021EBMF by setting $K=40$. For EBMR–sep, we discretize $[0,1]$ on a $K=10$ grid and parameterize $\textrm{g}_{\mathrm{sep}}$ as a two–layer ReLU network with 20 hidden units per layer. We train $\textrm{g}_{\textrm{sep}}$ with Adam (learning rate $0.01$) on the variational objective in Algorithm˜2 over 500 iterations with 50 gradient steps per iteration. The posterior sampling step uses NUTS with 300 warm–up and 300 retained samples (target acceptance probability $0.8$). Each EBMR–sep run requires approximately 5 minutes on a GPU; a full noise–level sweep takes two hours on two GPUs.
Figure˜4 shows the results. EBMR–sep attains uniformly lower R–MSE than both NPMLE and EBMF, with the performance gap increasing under higher noise. The intuitive reason is that the separately exchangeable prior captures gene-level and patient-level sharing of information in a nonlinear way, whereas the other two methods either ignore the row/column-exchangeable structure (NPMLE) or assume a low-rank structure (EBMF).
Figure 4: Breast cancer data denoising: EBMR with a separately exchangeable prior achieves the lowest relative MSE across varying noise levels. Scatterplots represent median relative MSE over 10 replicates; bands show the interquartile range (25–75%). There is a “blessing of noise” phenomenon: as the noise level increases, all empirical Bayes methods achieve lower relative MSE.
5.2 Denoising and Interpolating Air Quality Data in NYC
We apply empirical Bayes spatial regression (SpatialEB) to a dataset on fine particulate matter $\mathrm{PM}_{2.5}$ (in $\mu\mathrm{g}/\mathrm{m}^{3}$) in New York City. The data comprise hourly readings from 16 monitoring stations obtained from the NYC Department of Health and Mental Hygiene’s Real-Time Air Quality portal. [^2] https://a816-dohbesp.nyc.gov/IndicatorPublic/data-features/realtime-air-quality/ We focus on four sites—Herald Square, Chinatown, the Lower East Side, and Queens North—using records from January 2019 to May 2025.333We select these four sites because they have the most complete records; for most other sites (including those in Brooklyn, Staten Island, and the Bronx), more than $60\%$ of weekly observations are missing. To capture medium-term dynamics, we aggregate to weekly averages $x_{s,t}$ at site $s$ and week $t$ and assign sampling precisionconsis $\tau_{s,t}$ as the inverse within-week variances.
The weekly series contain gaps (outages/maintenance), occasional spikes, and heterogeneous noise. Our goals are to (i) denoise and interpolate the weekly average data for all four sites, (ii) estimate the latent weekly $\mathrm{PM}_{2.5}$ levels with uncertainty quantification, and (iii) characterize the spatio–temporal dependence structure via the fitted spectral density.
We model the data using the spatial regression model
$x_{s,t}\mid z_{s,t},\beta\sim\mathcal{N}\bigl(z_{s,t}+\beta^{\top}a_{s},\tau_{s,t}^{-1}\bigr),$
where $z_{s,t}$ denotes the latent weekly mean and $a_{s}$ is a site-specific covariate vector including an intercept and site indicators. The regression component $\beta^{\top}a_{s}$ accounts for persistent location effects across monitoring sites. We fit the spatial regression model with $K=10$ Gaussian–mixture components for the spectral density, then apply the procedure in Appendix˜E to obtain posterior means $\hat{z}_{s,t}$, regression coefficients $\hat{\beta}$, $95\%$ credible bands for $\{z_{s,t}\}$, and kriged imputations for missing weeks. We also report site-adjusted means $\hat{m}_{s,t}=\hat{z}_{s,t}+\hat{\beta}^{\top}a_{s}$ with credible bands.
When we parameterize the spectral density, $\omega=(\omega_{\text{time}},\omega_{\text{space}})$ denotes temporal and spatial frequencies. The model learns a spectral density $\psi(\omega)$; by Bochner’s theorem this induces a stationary covariance $k(\Delta)$ in the $(t,\text{space})$ domain. Peaks in $\psi(\omega)$ correspond to dominant temporal cycles and spatial scales of variation in $z_{s,t}$.
We assess the imputation accuracy of EB spatial regression against three Gaussian–process regressors on the same site–week features as our EB model, standardizing inputs and using per–observation noise variances $\alpha_{i}=\tau_{i}^{-1}$ from within-week precision. The models differ only by kernel—RBF (squared–exponential, smooth), Matérn $(\nu=3/2)$ (rougher trajectories), and Rational Quadratic (a scale–mixture kernel capturing multi–scale variability)—each with a small white-noise nugget ($10^{-3}$). Kernel hyperparameters are learned by maximizing the marginal likelihood. On held–out data, we report the MSE of the imputed series against the true weekly averages. Figure˜6 shows that EB spatial regression achieves substantially lower MSE than all GP baselines because the fitted spectral density adaptively captures the bimodal spatio–temporal variation. By comparing the four methods, we see that fitting a more flexible prior family that respects the shift invariance yields performance gain in imputation.
Figure˜5 shows empirical Bayes posterior means with $95\%$ credible bands. The posterior means shows the shrinkage and pooling effects of EB spatial regression: extreme weekly averages (e.g., mid-2023) are pulled toward the mean. Where observations are dense, the posterior uncertainty around the denoised weekly averages is narrow; where data are missing, it is wider. Within Manhattan (Herald Square, Chinatown, Lower East Side), the model borrows heavily across sites—for example, when Chinatown is entirely missing from mid-2021 to early-2022, the model interpolates from Herald Square and the Lower East Side to produce tight bands; similarly, during mid-2023 to early-2024, the model imputes the Lower East Side using the other two Manhattan sites, which recovers a decreasing, cyclical trend. By contrast, Queens North contributes less to interpolating Manhattan gaps (e.g., mid-2022 to early-2023). The reason for this different shrinking behavior is that the EB model finds two spatio–temporal regimes–one for Manhattan and one for northern Queens; see Figure˜9 in §˜G.3.
Figure 5: Denoised and interpolated weekly $\mathrm{PM}_{2.5}$ in New York City using EB spatial regression. Posterior means with 95% credible bands for three monitoring sites; gray points are observed weekly averages. The model smooths noisy measurements and imputes missing weeks.
Figure 6: Empirical Bayes spatial regression achieves substantially lower MSE than Gaussian–process baselines. Boxplots show test MSE (median and quantiles) across 50 random masks that hide 20% of observed entries in the NYC air–quality panel. Methods: EB spatial regression with $K=10$ Gaussian–mixture components for the spectral density, and GPR with RBF, Matérn ($\nu=3/2$), and Rational–Quadratic kernels.
6 Discussion
We have presented Bayesian empirical Bayes, a general empirical Bayes procedure to obtain accurate simultaneous posterior inference of latent variables when the prior respects a general probabilistic symmetry. Our approach enables the full power of empirical Bayes to be applied to problems with structured data, beyond exchangeable sequences. In this way, it expands the scope of application for EB substantially. Using Bayesian EB, we derived new EB models and inference algorithms for three settings: array data with separate and joint exchangeability, spatial data with shift invariance, and covariate-augmented data. We demonstrated the competitive performance of Bayesian EB over classical empirical Bayes.
There are several avenues for further research.
First, we have focused here on the derivation and methods behind Bayesian EB. Its theoretical analysis remains an open area. Natural theoretical questions include the identifiability, consistency, and sample complexity of learning the invariant variable g via maximum marginal likelihood and more broadly, estimation under symmetry constraints.
Second, in this paper, we studied symmetry groups based on arrays and spatio-temporal constraints. In many scientific domains, however, the relevant symmetry groups are more complex, such as those arising from physical constraints (e.g., molecular or spacetime symmetries). Another interesting direction is to leverage geometric deep learning (Bronstein2021) to design new BEB methods tailored to such applications.
Finally, while our initial evidence suggests that Bayesian EB remains competitive under heavy-tailed noise (see Figure˜8 in §˜G.2), it would be valuable to develop sensitivity analysis for these new methods under other types of model misspecification. Sensitivity analysis for empirical Bayes has been studied in Berger1986; Buta2011; Doss2024.
Appendix A Conditions on Groups
We provide the regularity conditions on the group of transformations $\Phi$ acting on the space $\mathcal{Z}^{\Omega}$. For $\phi_{1},\phi_{2}\in\Phi$, we write the composition $\phi_{1}\phi_{2}(\cdot):=\phi_{1}\bigl(\phi_{2}(\cdot)\bigr)$. If $\phi\in\Phi$ and $A\subseteq\Phi$, we write $\phi A:=\{\phi\psi\mid\psi\in A\}$. If $K,A\subseteq\Phi$, we write $KA:=\{\phi\psi\mid\phi\in K,\psi\in A\}$.
We equip $\Phi$ with the Borel $\sigma$-algebra $\mathcal{B}(\Phi)$. For every locally compact, second countable, and Hausdorff (lcscH) group, there exists a $\sigma$-finite measure $\lvert\cdot\rvert$ such that
$\bigl\lvert\phi^{-1}A\bigr\rvert=\lvert A\rvert,\quad\text{for all }\phi\in\Phi,\ A\in\mathcal{B}(\Phi),$
called a Haar measure (Folland2016). The Haar measure is unique up to a positive scaling factor. If $A\subset\Phi$ is compact, then $\lvert A\rvert$ is finite.
A Følner sequence in $\Phi$ is a sequence of compact sets $\bm{A}_{1},\bm{A}_{2},\dots$ such that, for every compact subset $K\subset\Phi$,
$\frac{\lvert K\bm{A}_{n}\cap\bm{A}_{n}\rvert}{\lvert\bm{A}_{n}\rvert}\xrightarrow[n\to\infty]{}1.$
If $\Phi$ is discrete, its compact sets are finite, and hence the condition specializes to the familiar finite Følner property. An lcscH group is called amenable if it admits a Følner sequence.
A Følner sequence $\{\bm{A}_{n}\}$ is tempered if there is some constant $c>0$ such that, for all $n\in\mathbb{N}$,
$\bigl\lvert\bigcup_{k<n}\bm{A}_{k}^{-1}\bm{A}_{n}\bigr\rvert\leq c\bigl\lvert\bm{A}_{n}\bigr\rvert$.
Not every Følner sequence is tempered, but every lcscH group that has a Følner sequence also has a tempered one (Lindenstrauss2001, Proposition 1.4).
Finally, we give the definition of a "nice" group, which matches the standard definition in the literature of probabilistic symmetry (Austern2022).
Definition 2 (Nice group).
We call $\Phi$ a nice group if
  1. (i)
    $\Phi$ is a locally compact, second countable, Hausdorff (lcscH) topological group, and the group operations are continuous.
  2. (ii)
    $\Phi$ is amenable, i.e., it admits a Følner sequence $\left(\bm{A}_{n}\right)_{n\in\mathbb{N}}$.
Appendix B Empirical Bayes Beyond Index Transformations
In §˜2, we assumed $\Phi$ to be a group of transformations acting on the index space $\Omega$. There are many interesting groups beyond index transformations, for example the rotation group or the Heisenberg group. We can develop Bayesian empirical Bayes methods for these groups under an equivariant assumption on the likelihood.
Let $\bm{x}\in\mathcal{X}^{\Omega}$ be observed variables and $\bm{z}\in\mathcal{X}^{\Omega}$ local latent variables. A general EB setup is as follows:
$\bm{z}\sim\mathrm{p}(\bm{z}),\quad\bm{x}\sim\mathrm{p}(\bm{x}\mid\bm{z}).$
The likelihood $\mathrm{p}(\bm{x}\mid\cdot)$ is assumed to be $\Phi$-equivariant in the sense of Definition˜4. Specifically,
$\mathrm{p}(\bm{x}\mid\phi(\bm{z}))=\mathrm{p}\left(\phi^{-1}(\bm{x})\mid\bm{z}\right),\quad\forall\phi\in\Phi.$
As a special case, the product likelihood $\mathrm{p}(\bm{x}\mid\bm{z})=\prod_{\omega\in\Omega}\mathrm{p}(x_{\omega}\mid z_{\omega})$ is equivariant with respect to any group of index transformations. The Gaussian likelihood is equivariant to rotations.
By Assumption˜1 and Assumption˜2, there exists a true prior $\mathrm{p}^{\star}(\bm{z})$ that is $\Phi$-invariant. If $\Phi$ is a nice group in the sense of Definition˜2, the ergodic representation theorem holds for $\Phi$. Thus, there exists an invariant variable g such that the population distribution is
$\textrm{g}\sim\mu^{\star},\quad\bm{z}\sim\mathrm{p}(\bm{z}\mid\textrm{g}),\quad\bm{x}\sim\mathrm{p}(\bm{x}\mid\bm{z}).$ (44)
The goal is to compute the posterior $\mathrm{p}^{\star}(\bm{z}\mid\bm{x})$, given by
$\mathrm{p}^{\star}(\bm{z}\mid\bm{x})\propto\int_{\mathcal{G}}\mathrm{p}(\bm{x}\mid\bm{z})\,\mathrm{p}(\bm{z}\mid\textrm{g})\,\mathrm{d}\mu^{\star}(\textrm{g}).$
[…]
Proposition 3.
The population posterior $\mathrm{p}^{\star}(\bm{z}\mid\bm{x})$ is a $\Phi$–equivariant conditional distribution, in the sense that for any $\phi\in\Phi$,
$\mathrm{p}^{\star}\!\left(\phi^{-1}(\bm{z})\mid\bm{x}\right)=\mathrm{p}^{\star}\!\left(\bm{z}\mid\phi(\bm{x})\right).$
[…]
This result shows that transforming the structured variable $\bm{z}$ by $\phi$ under the posterior is equivalent to computing the posterior of $\bm{z}$ after applying the transformation $\phi$ to the observed variables $\bm{x}$.
Proof.
We first consider the population posterior $\mathrm{p}^{\star}(\bm{z}\mid\bm{x})$. For any $\phi\in\Phi$,
$\mathrm{p}^{\star}\!\left(\phi(\bm{z})\mid\bm{x}\right)$ $=\int_{\mathcal{G}}\mathrm{p}\!\left(\bm{x}\mid\phi(\bm{z})\right)\,\mathrm{p}\!\left(\phi(\bm{z})\mid\textrm{g}\right)\,\mathrm{d}\mu^{\star}(\textrm{g})$
$=\int_{\mathcal{G}}\mathrm{p}\!\left(\bm{x}\mid\phi(\bm{z})\right)\,\mathrm{p}\!\left(\bm{z}\mid\textrm{g}\right)\,\mathrm{d}\mu^{\star}(\textrm{g})$
$=\int_{\mathcal{G}}\mathrm{p}\!\left(\phi^{-1}(\bm{x})\mid\bm{z}\right)\,\mathrm{p}\!\left(\bm{z}\mid\textrm{g}\right)\,\mathrm{d}\mu^{\star}(\textrm{g})=\mathrm{p}^{\star}\!\left(\bm{z}\mid\phi^{-1}(\bm{x})\right).$
[…]
The result suggests that “transform then condition” and “condition then transform” commute for the posterior. The population posterior mean $\mathbb{E}\left[\bm{z}\mid\bm{x}\right]:=\int_{\mathcal{X}^{\Omega}}\bm{z}\,\mathrm{p}^{\star}(\bm{z}\mid\bm{x})\,\mathrm{d}\bm{z}$ therefore satisfies $\mathbb{E}\left[\phi(\bm{z})\mid\bm{x}\right]=\mathbb{E}\left[\bm{z}\mid\phi(\bm{x})\right]$ for any $\phi\in\Phi$ when $\Phi$ is a group of isometries. This suggests that, to learn the population Bayes estimator of transformed latent variables, it suffices to compute the posterior on transformed datasets. This observation connects empirical Bayes point estimation (Jiang2009) to the common machine learning practice of enforcing symmetry via data augmentation (gonzalez2018galaxy; lyle2019; chen2020contrastive).
Appendix C Support Results
Definition 3.
A space $\Omega$ is exhaustible by compact sets if, for every open set $U\subseteq\Omega$, there exists an increasing sequence of compact subsets $S_{1}\subseteq S_{2}\subseteq\cdots$ such that
$U=\bigcup_{n=1}^{\infty}S_{n},$
and, for every $n$, $S_{n}$ is contained in the interior of $S_{n+1}$.
Definition 4 ($\Phi$-equivariance).
A function $t:\mathcal{X}^{\Omega}\to\mathcal{X}^{\Omega}$ is $\Phi$-equivariant if
$\phi(t(\bm{x}))=t(\phi(\bm{x})),\quad\text{for all }\phi\in\Phi.$ (45)
Intuitively, a function $t$ is $\Phi$-equivariant if every action of $\Phi$ “commutes” with $t$. An equivariant function $t$ preserves $\Phi$-invariance, as $\phi(\bm{x})\overset{d}{=}\bm{x}$ implies
$\phi(t(\bm{x}))=t(\phi(\bm{x}))\overset{d}{=}t(\bm{x}).$
Appendix D Algorithms for Empirical Bayes Matrix Recovery
This appendix details the inference algorithms for the empirical Bayes matrix–recovery methods introduced in §˜3. §˜D.1 presents a variational procedure for matrix recovery under a separately exchangeable prior (EBMR–sep; Algorithm˜2). §˜D.2 develops the inference algorithm for EBMR with a jointly exchangeable prior (EBMR–joint; Algorithm˜3), and §˜D.3.1 provides the algorithm for covariate-assisted EB (CAEB) which is analogous to the separately exchangeable case.
For all three algorithms, we use variational inference (Blei2017) to approximate the nonparametric maximum marginal likelihood.
Variational MMLE. Let $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{X}^{S_{n}})$ be a family of distributions. By the Gibbs variational principle (VanHandel2014, Lemma 4.10), one can lower bound the marginal likelihood by
$\log\mathrm{p}\bigl(\bm{x}_{n}\mid\textrm{g}\bigr)\geq-\ell(\bm{x}_{n};\textrm{g}):=\sup_{\textrm{q}\in\mathcal{Q}}\mathbb{E}_{\textrm{q}(\bm{z}_{n})}\left[\log\mathrm{p}(\bm{x}_{n}\mid\bm{z}_{n})+\log\mathrm{p}(\bm{z}_{n}\mid\textrm{g})-\log\textrm{q}(\bm{z}_{n})\right],$ (46)
where $\ell(\bm{x}_{n};\textrm{g})$ is the maximum evidence lower bound (ELBO) over the family $\mathcal{Q}$.
D.1 EBMR with a Separately Exchangeable Prior
To approximate the intractable marginal likelihood in Eq.˜35, we posit the following structured variational family
$\mathcal{Q}_{\textrm{sep}}=\left\{\textrm{q}\in\mathcal{P}\left([0,1]^{n+p+np}\right):\textrm{q}\left(\bm{u}_{n,p}\right)=\prod_{i=1}^{n}\textrm{q}_{\bm{u},i}(u_{i})\,\prod_{j=1}^{p}\textrm{q}_{\bm{v},j}(v_{j})\,\prod_{i,j}\textrm{q}_{i,j}(u_{ij}\mid u_{i},v_{j})\right\}.$ (47)
Figure˜7 motivates $\mathcal{Q}_{\textrm{sep}}$. In the exact posterior $\mathrm{p}(\bm{u}_{n,p}\mid\hat{\textrm{g}}_{\textrm{sep}},\bm{x}_{n,p})$, the variables $(u_{i})_{i\in[n]}$ and $(v_{j})_{j\in[p]}$ form a dense clique, while each $u_{ij}$ connects only to $(u_{i},v_{j})$. Our variational family replaces the dense clique with a product measure over $(u_{i})$ and $(v_{j})$ while preserving the local $u_{i}$$u_{ij}$$v_{j}$ chains.
[…]
Figure 7: Variational approximation. Left: exact posterior with a dense clique among $(u_{i})$ and $(v_{j})$ and local connections to $u_{ij}$. Right: variational family $\textrm{q}\in\mathcal{Q}_{\textrm{sep}}$ factorizes across $(u_{i})$ and $(v_{j})$ while preserving the local chain $u_{i}$$u_{ij}$$v_{j}$.
Maximizing the ELBO over $\mathcal{Q}_{\textrm{sep}}$ yields a lower bound of the marginal likelihood,
$\log\mathrm{p}\bigl(\bm{x}_{n,p}\mid\textrm{g}_{\textrm{sep}}\bigr)\geq\sup_{\textrm{q}\in\mathcal{Q}_{\textrm{sep}}}\operatorname{ELBO}\left(\textrm{q},\textrm{g}_{\textrm{sep}}\right):=\sup_{\textrm{q}\in\mathcal{Q}_{\textrm{sep}}}\mathbb{E}_{\textrm{q}(\bm{u}_{n,p})}\left[\log\mathrm{p}\bigl(\bm{x}_{n,p}\mid\textrm{g}_{\textrm{sep}},\bm{u}_{n,p}\bigr)-\log\textrm{q}(\bm{u}_{n,p})\right].$
Let $\textrm{q}_{\bm{u}}=\bigotimes_{i=1}^{n}\textrm{q}_{\bm{u},i}$ and $\textrm{q}_{\bm{v}}=\bigotimes_{j=1}^{p}\textrm{q}_{\bm{v},j}$. We estimate $\textrm{g}_{\textrm{sep}}$ by minimizing the variational loss
$\hat{\textrm{g}}_{\textrm{sep}}\in\arg\min_{\textrm{g}_{\textrm{sep}}\in\mathcal{G}}\ell(\bm{x}_{n,p};\textrm{g}_{\textrm{sep}})\coloneqq-\sup_{\textrm{q}\in\mathcal{Q}_{\textrm{sep}}}\textsc{ELBO}(\textrm{q},\textrm{g}_{\textrm{sep}}).$ (48)
A direct calculation (see §˜F.2) shows that
$\ell\left(\bm{x}_{n,p};\textrm{g}_{\textrm{sep}}\right)$ $=\inf_{\textrm{q}_{\bm{u}}\in\mathcal{P}([0,1])^{\otimes n},\textrm{q}_{\bm{v}}\in\mathcal{P}([0,1])^{\otimes p}}\Bigg\{-\sum_{i=1}^{n}\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})\,\textrm{q}_{\bm{v},j}(v_{j})}\left[\operatorname{LSE}\left(x_{ij},\tau_{ij},u_{i},v_{j},\textrm{g}_{\textrm{sep}}\right)\right]$ (49)
$\hskip 56.9055pt+\sum_{i=1}^{n}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})}\left[\log\textrm{q}_{\bm{u},i}(u_{i})\right]+\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{v},j}(v_{j})}\left[\log\textrm{q}_{\bm{v},j}(v_{j})\right]\Bigg\},$
where $\operatorname{LSE}\left(X,\tau,U,V,\textrm{g}_{\textrm{sep}}\right):=\log\int_{0}^{1}\exp\left(-\frac{\tau}{2}\left(X-\textrm{g}_{\textrm{sep}}(U,V,u)\right)^{2}\right)\,du$.
Alternating optimization.
To fit the variational distributions, we discretize $\textrm{q}_{\bm{u},i}$ and $\textrm{q}_{\bm{v},j}$ on the (K+1)(K+1)–point grid $u_{k}:=k/K$ for $k=0,\ldots,K$ over $[0,1]$:
$\textrm{q}_{\bm{u},i}=\sum_{k=0}^{K}w_{\bm{u},ik}\,\delta_{u_{k}},\quad\textrm{q}_{\bm{v},j}=\sum_{k=0}^{K}w_{\bm{v},jk}\,\delta_{u_{k}},\quad\text{where }u_{k}=k/K,\ \sum_{k=0}^{K}w_{\bullet,\cdot k}=1.$ (50)
Following Eq.˜49, the empirical Bayes estimate minimizes the negative ELBO over the network and grid weights:
$\begin{split}\left(\hat{\textrm{g}}_{\textrm{sep}},\hat{\bm{w}}_{\bm{u}},\hat{\bm{w}}_{\bm{v}}\right)\in\arg\min_{\textrm{g}_{\textrm{sep}}\in\mathcal{G}_{\textrm{NN}},\ \bm{w}_{\bm{u}},\bm{w}_{\bm{v}}\in\Delta_{K+1}}\Bigg\{&-\sum_{i=1}^{n}\sum_{j=1}^{p}\sum_{k=0}^{K}\sum_{k^{\prime}=0}^{K}w_{\bm{u},ik}\,w_{\bm{v},jk^{\prime}}\,\widehat{\operatorname{LSE}}\bigl(x_{ij},\tau_{ij},u_{k},u_{k^{\prime}},\textrm{g}_{\textrm{sep}}\bigr)\\ &\quad+\sum_{i=1}^{n}\sum_{k=0}^{K}w_{\bm{u},ik}\log w_{\bm{u},ik}+\sum_{j=1}^{p}\sum_{k=0}^{K}w_{\bm{v},jk}\log w_{\bm{v},jk}\Bigg\},\end{split}$ (51)
where $\widehat{\operatorname{LSE}}\bigl(x,\tau,u,v,\textrm{g}_{\textrm{sep}}\bigr)\coloneqq\log\sum_{k=0}^{K}\exp\!\left\{-\tfrac{\tau}{2}\bigl(x-\textrm{g}_{\textrm{sep}}(u,v,k/K)\bigr)^{2}\right\}-\log(K+1).$
We minimize (51) by applying a block–coordinate descent algorithm: (i) update the row weights $\{\bm{w}_{\bm{u},i}\}$, (ii) update the column weights $\{\bm{w}_{\bm{v},j}\}$, (iii) update the network parameters of $\textrm{g}_{\textrm{sep}}$ by stochastic gradient descent (Adam; Kingma2014Adam). Algorithm˜2 provides the explicit updates. The weight updates can be parallelized over $i$ and $j$. Each iteration costs $O(npK^{2})$ time and $O((n+p)K)$ memory.
Input: Data matrix $\bm{x}_{n,p}$, precision matrix $\bm{\tau}_{n,p}$, neural network class $\mathcal{G}_{\textrm{NN}}$, grid size $K$, tuning parameters (epochs, learning rate)
Initialize: neural network parameters; probability weights $\bm{w}_{\bm{u},1},\ldots,\bm{w}_{\bm{u},n},\ \bm{w}_{\bm{v},1},\ldots,\bm{w}_{\bm{v},p}\in\Delta_{K+1}$.
[2pt] repeat
 for $i=1$ to $n$ (parallelizable) do
$\rho_{\bm{u},ik}$ $\leftarrow\sum_{j=1}^{p}\sum_{k_{2}=0}^{K}w_{\bm{v},jk_{2}}\,\widehat{\operatorname{LSE}}(x_{ij},\tau_{ij},u_{k},u_{k_{2}},\textrm{g}_{\mathrm{sep}}),$ $k=0,\ldots,K,$
$\bm{w}_{\bm{u},i}$ $\leftarrow\textsc{softmax}(\rho_{\bm{u},i0},\ldots,\rho_{\bm{u},iK}).$
   end for
 for $j=1$ to $p$ (parallelizable) do
$\rho_{\bm{v},jk}$ $\leftarrow\sum_{i=1}^{n}\sum_{k_{1}=0}^{K}w_{\bm{u},ik_{1}}\,\widehat{\operatorname{LSE}}(x_{ij},\tau_{ij},u_{k_{1}},u_{k},\textrm{g}_{\mathrm{sep}}),$ $k=0,\ldots,K,$
$\bm{w}_{\bm{v},j}$ $\leftarrow\textsc{softmax}(\rho_{\bm{v},j0},\ldots,\rho_{\bm{v},jK}).$
   end for
  Update the neural network parameters by stochastic gradient descent:
$\textrm{g}_{\mathrm{sep}}\leftarrow\arg\max_{\textrm{g}_{\mathrm{sep}}\in\mathcal{G}_{\textrm{NN}}}\sum_{i=1}^{n}\sum_{j=1}^{p}\sum_{k_{1}=0}^{K}\sum_{k_{2}=0}^{K}w_{\bm{u},ik_{1}}w_{\bm{v},jk_{2}}\widehat{\operatorname{LSE}}\bigl(x_{ij},\tau_{ij},u_{k_{1}},u_{k_{2}},\textrm{g}_{\mathrm{sep}}\bigr).$
until convergence;
Output: Fitted weights $\bm{w}_{\bm{u},1},\ldots,\bm{w}_{\bm{u},n},\ \bm{w}_{\bm{v},1},\ldots,\bm{w}_{\bm{v},p}$ and network $\textrm{g}_{\mathrm{sep}}$.
Algorithm 2 EBMR with a Separately Exchangeable Prior (EBMR–sep)
Proposition 4.
Algorithm˜2 performs block–coordinate ascent on the variational inference problem (51).
Approximate posterior computation.
The last step of Algorithm˜1 is to sample from the empirical Bayes posterior
[…]
We take the default sampler to be the No-U-Turn Sampler (NUTS) (Hoffman2014), an adaptive variant of Hamiltonian Monte Carlo (Neal2010) implemented in the Pyro probabilistic programming package (Bingham2019). NUTS automatically tunes the step size and trajectory length parameters. We run NUTS for a fixed number of iterations that includes an initial warmup phase and thinning. When the full MCMC fails to converge, an alternative is to draw surrogate samples from variational approximation $\hat{\textrm{q}}(\bm{z}_{n})$ obtained in Algorithm˜2. This requires using the fitted variational distributions from Algorithm˜2 to approximately sample from the posterior $\mathrm{p}\left(\bm{z}_{n,p}\mid\hat{\textrm{g}}_{\textrm{sep}},\bm{x}_{n,p}\right)$. Suppose that Algorithm˜2 returns $\hat{\textrm{q}}_{\bm{u},i}=\sum_{k=0}^{K}\hat{w}_{\bm{u},ik}\,\delta_{u_{k}}$ and $\hat{\textrm{q}}_{\bm{v},j}=\sum_{k=0}^{K}\hat{w}_{\bm{v},jk}\,\delta_{u_{k}}$. To sample $z_{ij}$, we first draw $u_{i}\sim\hat{\textrm{q}}_{\bm{u},i}$ and $v_{j}\sim\hat{\textrm{q}}_{\bm{v},j}$. Conditional on $(u_{i},v_{j})$, we draw $u_{ij}\in\{0,1/K,\ldots,1\}$ with probability
$\hat{\textrm{q}}_{ij}\big(u_{ij}=k/K\mid u_{i},v_{j},\bm{x}_{n,p}\big)=\frac{\exp\!\left\{-\tfrac{\tau_{ij}}{2}\big(x_{ij}-\hat{\textrm{g}}_{\textrm{sep}}(u_{i},v_{j},k/K)\big)^{2}\right\}}{\sum_{\ell=0}^{K}\exp\!\left\{-\tfrac{\tau_{ij}}{2}\big(x_{ij}-\hat{\textrm{g}}_{\textrm{sep}}(u_{i},v_{j},\ell/K)\big)^{2}\right\}},\quad k=0,\ldots,K.$
This yields a sample $z_{ij}=\hat{\textrm{g}}_{\textrm{sep}}(u_{i},v_{j},u_{ij})$ from the variational posterior.
D.2 EBMR with a Jointly Exchangeable Prior
We now develop empirical Bayes matrix recovery under a jointly exchangeable prior. Assume that Assumption˜1 and Assumption˜2 hold with the true prior being jointly exchangeable. By the Aldous–Hoover theorem, the population distribution for $\bm{x}$ has the form
$\textrm{g}_{\mathrm{joint}}\sim\mu^{\star},$ (53)
$u_{i},\,u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],$ (54)
$z_{ij}=\textrm{g}_{\mathrm{joint}}(u_{i},u_{j},u_{ij}),$ (55)
$x_{ij}\sim\mathcal{N}\!\left(z_{ij},\tau_{ij}^{-1}\right),\quad(i,j)\in\mathbb{N}^{2}.$ (56)
For finite population, we require the observed matrix to have the same number of rows and columns to apply joint exchangeability. For $n\geq 1$, let $\bm{x}_{n}=(x_{\omega})_{\omega\in[n]^{2}}$ denote an observed $n\times n$ matrix. By Proposition˜1, $\bm{x}_{n}$ follows the marginal distribution of (54)–(56) on the index set $S_{n}=[n]^{2}$.
Let $\bm{z}_{n}=(z_{\omega})_{\omega\in S_{n}}$ denote the corresponding local latent variables. For the invariant variable $\textrm{g}_{\mathrm{joint}}$, the conditional distribution $\mathrm{p}\left(\bm{z}_{n}\mid\textrm{g}_{\mathrm{joint}}\right)$ is given by (54)–(55). By Proposition˜1, the finite-sample generative process is described by the same equations (53)–(56) with $(i,j)\in[n]^{2}$.
D.2.1 Fitting procedure
Let $\bm{u}_{n}=(u_{i},u_{ij})_{(i,j)\in[n]^{2}}$ denote the set of row-specific and entry-specific latent variables. Algorithm˜1 estimates $\hat{\textrm{g}}_{\mathrm{joint}}$ by maximizing the marginal likelihood
$\log\mathrm{p}\!\left(\bm{x}_{n}\mid\textrm{g}_{\mathrm{joint}}\right):=\log\int_{[0,1]^{\,n+n^{2}}}\prod_{i=1}^{n}\prod_{j=1}^{n}\mathrm{p}\!\left(x_{ij}\mid\textrm{g}_{\mathrm{joint}}\!\left(u_{i},u_{j},u_{ij}\right)\right)\,\mathrm{d}\bm{u}_{n}.$ (57)
Directly maximizing the marginal likelihood over all functions $\textrm{g}_{\mathrm{joint}}:[0,1]^{3}\mapsto\mathbb{R}$ is intractable, due to the $(n+n^{2})$-dimensional integral. So we approximate the integral using variational inference.
D.2.2 Variational approximation
Consider the posterior $\mathrm{p}\bigl(\bm{u}_{n}\mid\bm{x}_{n},\textrm{g}_{\mathrm{joint}}\bigr)$ under model (54)–(56).
In the exact posterior $\mathrm{p}\bigl(\bm{u}_{n}\mid\bm{x}_{n},\textrm{g}_{\mathrm{joint}}\bigr)$, the variables $(u_{ij})_{i,j\in[n]}$ are conditionally independent given $(u_{i})_{i\in[n]}$. To capture the graphical structure locally, we posit the variational family
$\mathcal{Q}_{\mathrm{joint}}=\Bigl\{\textrm{q}:\ \textrm{q}(\bm{u}_{n})=\prod_{i=1}^{n}\textrm{q}_{\bm{u},i}(u_{i})\,\prod_{i,j}\textrm{q}_{i,j}(u_{ij}\mid u_{i},u_{j})\Bigr\}.$ (58)
Let $\textrm{q}_{\bm{u}}=\bigotimes_{i=1}^{n}\textrm{q}_{\bm{u},i}$. The variational method maximizes the evidence lower bound (ELBO) subject to $\textrm{q}\in\mathcal{Q}_{\mathrm{joint}}$.
Proposition 5.
Let $\mathcal{Q}$ be the variational family in (58). The variational loss for $\textrm{g}_{\mathrm{joint}}$ is
$\ell\bigl(\bm{x}_{n};\textrm{g}_{\mathrm{joint}}\bigr)=\inf_{\textrm{q}_{\bm{u}}\in\mathcal{P}([0,1])^{\otimes n}}\Bigg\{$ $-\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})\,\textrm{q}_{\bm{u},j}(u_{j})}\left[\operatorname{LSE}\!\left(x_{ij},\tau_{ij},u_{i},u_{j},\textrm{g}_{\mathrm{joint}}\right)\right]$ (59)
$\qquad\qquad+\sum_{i=1}^{n}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})}\left[\log\textrm{q}_{\bm{u},i}(u_{i})\right]\Bigg\},$
where
$\operatorname{LSE}\!\left(x_{ij},\tau_{ij},u_{i},u_{j},\textrm{g}_{\mathrm{joint}}\right):=\log\int_{0}^{1}\exp\!\left(-\frac{\tau_{ij}}{2}\left[x_{ij}-\textrm{g}_{\mathrm{joint}}(u_{i},u_{j},u)\right]^{2}\right)\!\mathrm{d}u.$
We parametrize $\textrm{g}_{\mathrm{joint}}$ with a neural network class $\mathcal{G}_{\mathrm{NN}}$ and fit it by minimizing the variational loss:
$\hat{\textrm{g}}_{\mathrm{joint}}\in\arg\min_{\textrm{g}_{\mathrm{joint}}\in\mathcal{G}_{\mathrm{NN}}}\ell\bigl(\bm{x}_{n};\textrm{g}_{\mathrm{joint}}\bigr).$ (60)
Since $\textrm{q}_{\bm{u},i}\in\mathcal{P}([0,1])$, we discretize it as
$\textrm{q}_{\bm{u},i}=\sum_{k=0}^{K}w_{ik}\,\delta_{u_{k}},\qquad u_{k}=\frac{k}{K},\qquad\sum_{k=0}^{K}w_{ik}=1.$
Algorithm˜3 implements EBMR with a jointly exchangeable prior.
With the estimated function $\hat{\textrm{g}}_{\mathrm{joint}}$ from Algorithm˜3, we use the No–U–Turn sampler (NUTS) to sample from the posterior $\mathrm{p}\!\left(z_{ij}\mid\hat{\textrm{g}}_{\mathrm{joint}},\bm{x}_{n}\right)$ of the EBMR model, given by
$(u_{i},u_{j},u_{ij})\sim\mathrm{p}\!\left(u_{i},u_{j},u_{ij}\mid\bm{x}_{n},\hat{\textrm{g}}_{\mathrm{joint}}\right),\quad z_{ij}=\hat{\textrm{g}}_{\mathrm{joint}}(u_{i},u_{j},u_{ij}).$ (61)
Input: Data matrix $\bm{x}_{n}$, precision matrix $\tau_{n}=(\tau_{ij})_{i,j\in[n]}$, network class $\mathcal{G}_{\mathrm{NN}}$, grid size $K$, optimization hyperparameters (epochs, learning rate)
Initialize: network $\textrm{g}_{\mathrm{joint}}\in\mathcal{G}_{\mathrm{NN}}$; weights $\bm{w}_{1},\ldots,\bm{w}_{n}\in\Delta_{K+1}$.
[2pt]
repeat
 for $i=1$ to $n$ (parallelizable) do
$\rho_{ik}$ $\leftarrow\widehat{\operatorname{LSE}}\!\left(x_{ii},\tau_{ii},u_{k},u_{k},\textrm{g}_{\mathrm{joint}}\right)+\sum_{j\neq i}\sum_{k_{2}=0}^{K}w_{jk_{2}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k},u_{k_{2}},\textrm{g}_{\mathrm{joint}}\right),$ $k=0,\ldots,K,$
$\bm{w}_{i}$ $\leftarrow\textsc{softmax}\bigl(\rho_{i0},\ldots,\rho_{iK}\bigr).$
   end for
  Update the neural network parameters with stochastic gradient ascent:
$\textrm{g}_{\mathrm{joint}}\leftarrow\arg\max_{\textrm{g}\in\mathcal{G}_{\mathrm{NN}}}\sum_{i=1}^{n}\sum_{k=0}^{K}w_{ik}\!\left[\widehat{\operatorname{LSE}}\!\left(x_{ii},\tau_{ii},u_{k},u_{k},\textrm{g}\right)+\sum_{j\neq i}\sum_{k_{2}=0}^{K}w_{jk_{2}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k},u_{k_{2}},\textrm{g}\right)\right].$
until convergence;
Output: Fitted weights $\bm{w}_{1},\ldots,\bm{w}_{n}$ and network $\textrm{g}_{\mathrm{joint}}$.
Algorithm 3 EBMR with a Jointly Exchangeable Prior (EBMR–Joint)
Corollary 1.
Algorithm˜2 performs block–coordinate ascent on the variational inference problem (51).
D.3 Covariate-Assisted Empirical Bayes
We extend empirical Bayes matrix recovery with a separately exchangeable prior to incorporate covariates on rows and columns. Let $\Omega=\mathbb{N}^{2}$ and $\mathcal{X}=\mathbb{R}$. We observe a matrix $\bm{x}_{n,p}=(x_{ij})_{(i,j)\in[n]\times[p]}$ with row covariates $\bm{y}_{n}=(y_{i})_{i=1}^{n}$ with $y_{i}\in\mathcal{Y}$ and column covariates $\bm{a}_{p}=(a_{j})_{j=1}^{p}$ with $a_{j}\in\mathcal{A}$, generated via
$x_{ij}\sim\mathcal{N}(z_{ij},\tau_{ij}^{-1}),\qquad\tau_{ij}>0.$
The statistical goal is to recover the latent array $\bm{z}_{n,p}$. On the population level, we impose relative exchangeability with respect to the observed covariates: the infinite array $\bm{z}$ is said to be $\bm{y}$$\bm{a}$ exchangeable if its law is invariant under relabelings that preserve covariate values,
[…]
Under $\bm{y}$$\bm{a}$ exchangeability, an extension of Aldous–Hoover theorem yields a measurable representation (Crane2018): there exists a random function $\textrm{g}_{\textrm{rel}}:\mathcal{Y}\times\mathcal{A}\times[0,1]^{3}\to\mathbb{R}$ such that, for iid $u_{i},v_{j},u_{ij}\sim\textrm{unif}[0,1]$,
$u_{i},v_{j},u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\qquad z_{ij}=\textrm{g}_{\textrm{rel}}\bigl(y_{i},a_{j},u_{i},v_{j},u_{ij}\bigr).$
The covariate-assisted array model is
$u_{i},v_{j},u_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\quad z_{ij}=\textrm{g}_{\textrm{rel}}\left(y_{i},a_{j},u_{i},v_{j},u_{ij}\right),\quad x_{ij}\sim\mathcal{N}\left(z_{ij},\tau_{ij}^{-1}\right).$ (62)
We call the Bayesian empirical Bayes method for this model covariate-assisted empirical Bayes (CAEB). The estimation procedure mirrors EBMR with a separately exchangeable prior: we maximize a variational lower bound to the marginal likelihood over a structured family for $(u_{i},v_{j},u_{ij})$. We parameterize the invariant variable $\textrm{g}_{\textrm{rel}}$ as a neural network that takes $(y_{i},a_{j},u_{i},v_{j},u_{ij})$ as inputs. With $\hat{\textrm{g}}_{\textrm{rel}}$ in hand, we draw from the plug-in posterior $\mathrm{p}(\bm{z}_{n,p}\mid\hat{\textrm{g}}_{\textrm{rel}},\bm{x}_{n,p})$. In the next section, we derive the variational lower bound to learn the invariant variable $\textrm{g}_{\textrm{rel}}$.
Related approaches include the covariate-assisted EB method of Ignatiadis2019, which fits a normal-means sequence model with the prior’s mean and variance modeled as functions of covariates, and denault2025covariate, which develops covariate-assisted empirical Bayes matrix factorization. In the latter model, the loading and factor components each follow i.i.d. priors parameterized by row/column covariates.
D.3.1 Fitting procedure
Let $\bm{u}_{n,p}\coloneqq\bigl(u_{i},v_{j},u_{ij}\bigr)_{(i,j)\in[n]\times[p]}$. To follow Algorithm˜1, the inference steps of covariate-assisted EB involve
  1. 1.
    Compute $\hat{\textrm{g}}_{\textrm{rel}}$ by maximizing the marginal likelihood
    $\log\mathrm{p}\bigl(\bm{x}_{n,p}\mid\textrm{g}_{\textrm{rel}}\bigr)\coloneqq\log\int_{[0,1]^{n+p+np}}\exp\left\{-\sum_{i=1}^{n}\sum_{j=1}^{p}\frac{\tau_{ij}}{2}\bigl[x_{ij}-\textrm{g}_{\textrm{rel}}(y_{i},a_{j},u_{i},v_{j},u_{ij})\bigr]^{2}\right\}\,d\bm{u}_{n,p}.$ (63)
  2. 2.
    Sample $z_{ij}$ from the posterior $\mathrm{p}\left(z_{ij}\mid\hat{\textrm{g}}_{\textrm{rel}},\bm{x}_{n,p}\right)$ for $i\in[n]$ and $j\in[p]$.
As in EBMR–sep, the MMLE in (63) is intractable for large $n,p$. We therefore adopt the same structured variational family $\mathcal{Q}_{\textrm{sep}}$ from (47) for the latent variables $\bm{u}_{n,p}$ and discretize each variational marginal on a $K$-point grid over $[0,1]$. We parameterize $\textrm{g}_{\textrm{rel}}$ with a neural network class $\mathcal{G}_{\mathrm{NN}}$ and minimize the variational loss
$\hat{\textrm{g}}_{\textrm{rel}}\in\arg\min_{\textrm{g}_{\textrm{rel}}\in\mathcal{G}_{\mathrm{NN}}}\ell(\bm{x}_{n,p};\textrm{g}_{\textrm{rel}}),$
where, with discretized weights,
$\ell(\bm{x}_{n,p};\textrm{g}_{\textrm{rel}})=\;\inf_{\bm{w}_{\bm{u}},\bm{w}_{\bm{v}}\in\Delta_{K+1}}\Bigg\{$ $-\sum_{i=1}^{n}\sum_{j=1}^{p}\sum_{k=0}^{K}\sum_{k^{\prime}=0}^{K}\bm{w}_{\bm{u},ik}\,\bm{w}_{\bm{v},jk^{\prime}}\,\operatorname{LSE}\left(x_{ij},\tau_{ij},y_{i},a_{j},\tfrac{k}{K},\tfrac{k^{\prime}}{K},\textrm{g}_{\textrm{rel}}\right)$
$\quad+\sum_{i=1}^{n}\sum_{k=0}^{K}\bm{w}_{\bm{u},ik}\log\bm{w}_{\bm{u},ik}+\sum_{j=1}^{p}\sum_{k=0}^{K}\bm{w}_{\bm{v},jk}\log\bm{w}_{\bm{v},jk}\Bigg\},$
with
$\operatorname{LSE}(X,\tau,Y,A,U,V,\textrm{g}_{\textrm{rel}})\coloneqq\log\int_{0}^{1}\exp\left\{-\tfrac{\tau}{2}\bigl(X-\textrm{g}_{\textrm{rel}}(Y,A,U,V,u)\bigr)^{2}\right\}\,du.$
Weight updates are similar to those in EBMR–sep:
$\bm{w}_{\bm{u},ik}$ $\propto\exp\left(\sum_{j=1}^{p}\sum_{k^{\prime}=0}^{K}\bm{w}_{\bm{v},jk^{\prime}}\,\log\sum_{k^{\prime\prime}=0}^{K}\exp\left\{-\tfrac{\tau_{ij}}{2}\Big[x_{ij}-\textrm{g}_{\textrm{rel}}\bigl(y_{i},a_{j},\tfrac{k}{K},\tfrac{k^{\prime}}{K},\tfrac{k^{\prime\prime}}{K}\bigr)\Big]^{2}\right\}\right),$ (64)
$\bm{w}_{\bm{v},jk}$ $\propto\exp\left(\sum_{i=1}^{n}\sum_{k^{\prime}=0}^{K}\bm{w}_{\bm{u},ik^{\prime}}\,\log\sum_{k^{\prime\prime}=0}^{K}\exp\left\{-\tfrac{\tau_{ij}}{2}\Big[x_{ij}-\textrm{g}_{\textrm{rel}}\bigl(y_{i},a_{j},\tfrac{k^{\prime}}{K},\tfrac{k}{K},\tfrac{k^{\prime\prime}}{K}\bigr)\Big]^{2}\right\}\right),$ (65)
normalized to lie in $\Delta_{K+1}$. With fixed weights, update $\textrm{g}_{\textrm{rel}}$ by
$\hat{\textrm{g}}_{\textrm{rel}}\in\arg\max_{\textrm{g}_{\textrm{rel}}\in\mathcal{G}_{\textrm{NN}}}\sum_{i=1}^{n}\sum_{j=1}^{p}\sum_{k=0}^{K}\sum_{k^{\prime}=0}^{K}\bm{w}_{\bm{u},ik}\,\bm{w}_{\bm{v},jk^{\prime}}\,\log\sum_{k^{\prime\prime}=0}^{K}\exp\left\{-\tfrac{\tau_{ij}}{2}\Big[x_{ij}-\textrm{g}_{\textrm{rel}}\bigl(y_{i},a_{j},\tfrac{k}{K},\tfrac{k^{\prime}}{K},\tfrac{k^{\prime\prime}}{K}\bigr)\Big]^{2}\right\}.$ (66)
The algorithm fits $\hat{\textrm{g}}_{\textrm{rel}}$ by iterating steps (64)–(66) until convergence. With the estimate $\hat{\textrm{g}}_{\textrm{rel}}$, we sample from the posterior $\mathrm{p}\!\left(z_{ij}\mid\hat{\textrm{g}}_{\textrm{rel}},\bm{x}_{n,p}\right)$ of model (62), given by
$(u_{i},u_{j},u_{ij})\sim\mathrm{p}\!\left(u_{i},u_{j},u_{ij}\mid\hat{\textrm{g}}_{\textrm{rel}},\bm{x}_{n,p}\right),\quad z_{ij}=\hat{\textrm{g}}_{\textrm{rel}}(u_{i},u_{j},u_{ij}).$ (67)
Appendix E Algorithms for Empirical Bayes Spatial Regression
Recall that the invariant variable is the spectral density $\psi\in\mathcal{P}_{ac}(\mathbb{R}^{d})$. Following Wilson2013, we further parametrize $\psi$ as a Gaussian mixture,
$\psi_{\theta}(s)=\sum_{k=1}^{K}w_{k}\,\phi\!\left(s;\mu_{k},\Sigma_{k}\right),\qquad w_{k}\geq 0,\ \sum_{k=1}^{K}w_{k}=1,$ (68)
where each $\Sigma_{k}=\mathrm{diag}(\sigma_{k,1}^{2},\ldots,\sigma_{k,d}^{2})$ is diagonal. By Bochner’s theorem, Eq.˜68 induces the kernel
$k_{\theta}(\Delta)=\sum_{k=1}^{K}w_{k}\prod_{j=1}^{d}\cos\!\bigl(2\pi\,\mu_{k,j}\,\Delta_{j}\bigr)\,\exp\!\bigl(-2\pi^{2}\,\sigma_{k,j}^{2}\,\Delta_{j}^{2}\bigr),$ (69)
for lags $\Delta\in\mathbb{R}^{d}$. The invariant variable is thus $\theta=(w_{1:K},\mu_{1:K},\Sigma_{1:K})$, which governs the stationary prior $\mathrm{GP}\!\left(0,k_{\theta}\right)$ for the latent process.
With this parametrization, the empirical Bayes spatial regression model is
$\bm{z}\sim\mathrm{GP}(0,k_{\theta}),\qquad\beta\sim\pi(\beta)\propto 1,\qquad x_{\omega}\sim\mathcal{N}\!\bigl(\beta^{\top}a_{\omega}+z_{\omega},\ \tau_{\omega}^{-1}\bigr).$ (70)
Posterior Inference.
We estimate the kernel parameters $\theta$ by maximizing the marginal likelihood,
$\hat{\theta}\in\operatorname*{arg\,max}_{\theta\in\Theta}\log\int\Bigl[\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\mid z_{\omega},\beta)\Bigr]\,\mathrm{p}(\bm{z}_{n}\mid\theta)\,\pi(\beta)\;d\beta\,d\bm{z}_{n}.$ (71)
[…]
Let $S_{n}=(\omega_{1},\ldots,\omega_{n})$, $\bm{A}_{n}=[a_{\omega_{1}},\ldots,a_{\omega_{n}}]^{\top}\in\mathbb{R}^{n\times p}$, and $D_{\tau}=\mathrm{diag}(\tau_{\omega_{1}},\ldots,\tau_{\omega_{n}})$. With $\Sigma_{\theta}=[k_{\theta}(\omega_{i}-\omega_{j})]_{(i,j)\in[n]^{2}}$, define $S_{\theta}=\Sigma_{\theta}+D_{\tau}^{-1}$ and $W_{\theta}=S_{\theta}^{-1}$. Assuming $\bm{A}_{n}$ has full column rank and $\pi(\beta)\equiv 1$, the MMLE objective is
$\ell(\bm{x}_{n};\theta)=\log\det S_{\theta}+\log\!\bigl|\bm{A}_{n}^{\top}W_{\theta}\bm{A}_{n}\bigr|+\bm{x}_{n}^{\top}W_{\theta}\bm{x}_{n}-\bm{x}_{n}^{\top}W_{\theta}\bm{A}_{n}\hat{\beta}_{\theta},$ (72)
where $\hat{\beta}_{\theta}=(\bm{A}_{n}^{\top}W_{\theta}\bm{A}_{n})^{-1}\bm{A}_{n}^{\top}W_{\theta}\bm{x}_{n}$ is the weighted least-squares estimator.
We minimize (72) using stochastic gradient descent (Robbins1951). The computational bottleneck is to solve linear systems in $S_{\theta}$; we avoid forming $W_{\theta}$ explicitly by using a Cholesky factorization $S_{\theta}=LL^{\top}$. The resulting costs are $O(n^{3})$ time and $O(n^{2})$ space, which are tractable if the number of sites $n$ is in the thousands. Scalable alternatives are available, including Nyström approximations and variational low-rank Gaussian processes (hensman2013GP).
With $S_{\hat{\theta}}=\Sigma_{\hat{\theta}}+D_{\tau}^{-1}$, $W_{\hat{\theta}}=S_{\hat{\theta}}^{-1}$, and $\hat{\beta}_{\hat{\theta}}=(\bm{A}_{n}^{\top}W_{\hat{\theta}}\bm{A}_{n})^{-1}\bm{A}_{n}^{\top}W_{\hat{\theta}}\bm{x}_{n}$,
$\beta\mid\bm{x}_{n}\sim\mathcal{N}\!\left(\hat{\beta}_{\hat{\theta}},\,(\bm{A}_{n}^{\top}W_{\hat{\theta}}\bm{A}_{n})^{-1}\right),$ (73)
and, conditional on $\beta$,
$\bm{z}_{n}\mid\beta,\bm{x}_{n}\sim\mathcal{N}\!\left((\Sigma_{\hat{\theta}}^{-1}+D_{\tau})^{-1}D_{\tau}(\bm{x}_{n}-\bm{A}_{n}\beta),\,(\Sigma_{\hat{\theta}}^{-1}+D_{\tau})^{-1}\right).$ (74)
Posterior samples are obtained by Monte Carlo: sample $\beta^{(s)}$ from (73), then $\bm{z}_{n}^{(s)}$ from (74) with $\beta=\beta^{(s)}$.
Kriging.
To predict at new sites $S_{n}^{\prime}\supset S_{n}$, define $Q_{\theta}\in\mathbb{R}^{|S_{n}^{\prime}|\times|S_{n}^{\prime}|}$ by $Q_{\theta,ii}=[\Sigma_{\theta}^{-1}]_{ii}+\tau_{i}$ and $Q_{\theta,ij}=[\Sigma_{\theta}^{-1}]_{ij}$ for $i\neq j$, where $\tau_{i}$ denotes the noise precision at site $\omega_{i}$. Let $L\in\mathbb{R}^{|S_{n}^{\prime}|}$ have entries $L_{S_{n}}=D_{\tau}\bigl(\bm{x}_{S_{n}}-\bm{A}_{S_{n}}\beta\bigr)$ and zeros elsewhere. Then
$\bm{z}_{S_{n}^{\prime}}\mid\beta,\bm{x}_{n}\sim\mathcal{N}\!\left(Q_{\hat{\theta}}^{-1}L,\;Q_{\hat{\theta}}^{-1}\right).$ (75)
Appendix F Proofs
F.1 Proofs of §˜2
Proof of Proposition˜1 and Proposition˜2.
Let $S_{n}^{c}\coloneqq\Omega\setminus S_{n}$. The true distribution $\mathrm{p}^{\star}(\bm{x}_{n})$ is the $S_{n}$–marginal of $\mathrm{p}^{\star}(\bm{x})$:
$\mathrm{p}^{\star}(\bm{x}_{n})=\int\mathrm{p}^{\star}(\bm{x})\,\mathrm{d}\bm{x}_{S_{n}^{c}}.$
By Assumption˜1,
$\mathrm{p}^{\star}(\bm{x})=\int_{\mathcal{Z}^{\Omega}}\mathrm{p}^{\star}(\bm{z})\,\prod_{\omega\in\Omega}\mathrm{p}(x_{\omega}\mid z_{\omega})\,\mathrm{d}\bm{z}.$
Applying Fubini’s theorem to interchange the integrals over $\bm{z}$ and $\bm{x}_{S_{n}^{c}}$ gives
$\mathrm{p}^{\star}(\bm{x}_{n})$ $=\int_{\mathcal{X}^{S_{n}^{c}}}\int_{\mathcal{Z}^{\Omega}}\mathrm{p}^{\star}(\bm{z})\,\prod_{\omega\in\Omega}\mathrm{p}(x_{\omega}\mid z_{\omega})\,\mathrm{d}\bm{z}\,\mathrm{d}\bm{x}_{S_{n}^{c}}$
$=\int_{\mathcal{Z}^{\Omega}}\mathrm{p}^{\star}(\bm{z})\,\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\mid z_{\omega})\left(\int_{\mathcal{X}^{S_{n}^{c}}}\prod_{\omega\in S_{n}^{c}}\mathrm{p}(x_{\omega}\mid z_{\omega})\,\mathrm{d}\bm{x}_{S_{n}^{c}}\right)\mathrm{d}\bm{z}$
$=\int_{\mathcal{Z}^{\Omega}}\mathrm{p}^{\star}(\bm{z})\,\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\mid z_{\omega})\,\mathrm{d}\bm{z},$
which establishes Proposition˜1.
For Proposition˜2, we integrate the representation (10) over $\bm{x}_{S_{n}}$:
$\mathrm{p}^{\star}(\bm{x}_{n})$ $=\int_{\mathcal{Z}^{S_{n}}}\int_{\mathcal{G}}\!\left(\int_{\mathcal{Z}^{S_{n}^{c}}}\mathrm{p}(\bm{z}\mid\textrm{g})\,\mathrm{d}\bm{z}_{S_{n}^{c}}\right)\mathrm{d}\mu^{\star}(\textrm{g})\,\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\mid z_{\omega})\,\mathrm{d}\bm{z}_{S_{n}}$
$=\int_{\mathcal{Z}^{S_{n}}}\left(\int_{\mathcal{G}}\mathrm{p}(\bm{z}_{n}\mid\textrm{g})\,\mathrm{d}\mu^{\star}(\textrm{g})\right)\prod_{\omega\in S_{n}}\mathrm{p}(x_{\omega}\mid z_{\omega})\,\mathrm{d}\bm{z}_{S_{n}}.$
F.2 Proofs of the Algorithms
Proof of Eq.˜49.
For $\textrm{q}\in\mathcal{Q}_{\textrm{sep}}$ we can decompose the ELBO as
$\begin{split}\operatorname{ELBO}\left(\textrm{q},\textrm{g}_{\textrm{sep}}\right)=\,&\sum_{i=1}^{n}\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})\,\textrm{q}_{\bm{v},j}(v_{j})}\left[\mathbb{E}_{\textrm{q}_{ij}(u_{ij}\mid u_{i},v_{j})}\left[\log\mathrm{p}\!\left(x_{ij}\mid\textrm{g}_{\textrm{sep}}(u_{i},v_{j},u_{ij})\right)\right]\right]\\ &\quad-\sum_{i=1}^{n}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})}\left[\log\textrm{q}_{\bm{u},i}(u_{i})\right]-\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{v},j}(v_{j})}\left[\log\textrm{q}_{\bm{v},j}(v_{j})\right]\\ &\quad-\sum_{i=1}^{n}\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})\,\textrm{q}_{\bm{v},j}(v_{j})}\left[\mathbb{E}_{\textrm{q}_{ij}(u_{ij}\mid u_{i},v_{j})}\left[\log\textrm{q}_{ij}(u_{ij}\mid u_{i},v_{j})\right]\right].\end{split}$
Introduce the reference distribution
$\textrm{q}_{\mathrm{ref}}(u_{ij}\mid u_{i},v_{j})\propto\exp\!\left\{-\frac{\tau_{ij}}{2}\bigl(x_{ij}-\textrm{g}_{\textrm{sep}}(u_{i},v_{j},u_{ij})\bigr)^{2}\right\}.$
Then
$\operatorname{ELBO}\left(\textrm{q},\textrm{g}_{\textrm{sep}}\right)$ $=\sum_{i=1}^{n}\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})\,\textrm{q}_{\bm{v},j}(v_{j})}\left[\operatorname{LSE}(x_{ij},\tau_{ij},u_{i},v_{j},\textrm{g}_{\textrm{sep}})\right]$ (76)
$\quad-\sum_{i=1}^{n}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})}\left[\log\textrm{q}_{\bm{u},i}(u_{i})\right]-\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{v},j}(v_{j})}\left[\log\textrm{q}_{\bm{v},j}(v_{j})\right]$
$\quad-\sum_{i=1}^{n}\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})\,\textrm{q}_{\bm{v},j}(v_{j})}\left[D_{\textsc{KL}}\left(\textrm{q}_{ij}(\cdot\mid u_{i},v_{j})\parallel\textrm{q}_{\mathrm{ref}}(\cdot\mid u_{i},v_{j})\right)\right],$
where
$\operatorname{LSE}(X,\tau,U,V,\textrm{g}_{\textrm{sep}})\coloneqq\log\!\int_{0}^{1}\exp\!\left\{-\frac{\tau}{2}\bigl[X-\textrm{g}_{\textrm{sep}}(U,V,u)\bigr]^{2}\right\}\mathrm{d}u.$
Maximizing over $\textrm{q}_{ij}$ for fixed $\textrm{q}_{\bm{u}},\textrm{q}_{\bm{v}}$ gives
$\textrm{q}_{ij}^{\star}(u_{ij}\mid u_{i},v_{j})=\textrm{q}_{\mathrm{ref}}(u_{ij}\mid u_{i},v_{j}),$
and evaluating at $\textrm{q}^{\star}$ yields
$\operatorname{ELBO}\left(\textrm{q}^{\star},\textrm{g}_{\textrm{sep}}\right)$ $=\sum_{i=1}^{n}\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}^{\star}_{\bm{u},i}(u_{i})\,\textrm{q}^{\star}_{\bm{v},j}(v_{j})}\left[\operatorname{LSE}(x_{ij},\tau_{ij},u_{i},v_{j},\textrm{g}_{\textrm{sep}})\right]$
$\quad-\sum_{i=1}^{n}\mathbb{E}_{\textrm{q}^{\star}_{\bm{u},i}(u_{i})}\left[\log\textrm{q}^{\star}_{\bm{u},i}(u_{i})\right]-\sum_{j=1}^{p}\mathbb{E}_{\textrm{q}^{\star}_{\bm{v},j}(v_{j})}\left[\log\textrm{q}^{\star}_{\bm{v},j}(v_{j})\right].$
Proof of Proposition˜4.
Recall that $u_{k}=k/K$ for $k=0,\ldots,K$. The discretized form of (49) is
$\hat{\textrm{g}}_{\mathrm{sep}}\in\arg\max_{\textrm{g}_{\mathrm{sep}}\in\mathcal{G}_{\mathrm{NN}}}\ \sup_{\bm{w}\in\Delta_{K+1}^{\,n+p}}\Bigg\{$ $\sum_{i=1}^{n}\sum_{j=1}^{p}\sum_{k_{1}=0}^{K}\sum_{k_{2}=0}^{K}w_{\bm{u},ik_{1}}\,w_{\bm{v},jk_{2}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k_{1}},u_{k_{2}},\textrm{g}_{\mathrm{sep}}\right)$ (77)
$\quad-\sum_{i=1}^{n}\sum_{k=0}^{K}w_{\bm{u},ik}\log w_{\bm{u},ik}-\sum_{j=1}^{p}\sum_{k=0}^{K}w_{\bm{v},jk}\log w_{\bm{v},jk}\Bigg\},$
where $\bm{w}$ collects all weight vectors $\bm{w}_{\bm{u},1},\ldots,\bm{w}_{\bm{u},n},\bm{w}_{\bm{v},1},\ldots,\bm{w}_{\bm{v},p}$ and $\widehat{\operatorname{LSE}}$ is the discretized form of $\operatorname{LSE}$.
We fit $\bm{w}_{\bm{u}},\bm{w}_{\bm{v}}$ and $\textrm{g}_{\mathrm{sep}}$ by alternating updates. For $j\in[p]$ and $k=0,\ldots,K$,
$w_{\bm{v},jk}=\frac{\exp\!\left(\sum_{i=1}^{n}\sum_{k_{1}=0}^{K}w_{\bm{u},ik_{1}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k_{1}},u_{k},\textrm{g}_{\mathrm{sep}}\right)\right)}{\sum_{k^{\prime}=0}^{K}\exp\!\left(\sum_{i=1}^{n}\sum_{k_{1}=0}^{K}w_{\bm{u},ik_{1}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k_{1}},u_{k^{\prime}},\textrm{g}_{\mathrm{sep}}\right)\right)}.$ (78)
Similarly, for $i\in[n]$ and $k=0,\ldots,K$,
$w_{\bm{u},ik}=\frac{\exp\!\left(\sum_{j=1}^{p}\sum_{k_{2}=0}^{K}w_{\bm{v},jk_{2}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k},u_{k_{2}},\textrm{g}_{\mathrm{sep}}\right)\right)}{\sum_{k^{\prime}=0}^{K}\exp\!\left(\sum_{j=1}^{p}\sum_{k_{2}=0}^{K}w_{\bm{v},jk_{2}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k^{\prime}},u_{k_{2}},\textrm{g}_{\mathrm{sep}}\right)\right)}.$ (79)
Finally, we train the neural network $\textrm{g}_{\mathrm{sep}}$ by maximizing the weighted objective
$\textrm{g}_{\mathrm{sep}}\in\arg\max_{\textrm{g}_{\mathrm{sep}}\in\mathcal{G}_{\mathrm{NN}}}\sum_{i=1}^{n}\sum_{j=1}^{p}\sum_{k_{1}=0}^{K}\sum_{k_{2}=0}^{K}w_{\bm{u},ik_{1}}\,w_{\bm{v},jk_{2}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k_{1}},u_{k_{2}},\textrm{g}_{\mathrm{sep}}\right).$ (80)
Proof of Proposition˜5.
By the Gibbs variational principle,
$\log\mathrm{p}\!\left(\bm{x}_{n}\mid\textrm{g}_{\mathrm{joint}}\right)=\sup_{\textrm{q}\in\mathcal{P}([0,1]^{\,n+n^{2}})}\Bigg\{$ $\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbb{E}_{\textrm{q}(u_{i},u_{j},u_{ij})}\left[\log\mathrm{p}\!\left(x_{ij}\mid\textrm{g}_{\mathrm{joint}}(u_{i},u_{j},u_{ij})\right)\right]-\mathbb{E}_{\textrm{q}(\bm{u}_{n})}\left[\log\textrm{q}(\bm{u}_{n})\right]\Bigg\}.$ (81)
Leveraging the structure of $\textrm{q}\in\mathcal{Q}$ and rearranging yields
$\textsc{ELBO}=\sup_{\textrm{q}\in\mathcal{Q}}\Bigg\{-\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})\,\textrm{q}_{\bm{u},j}(u_{j})}\left[D_{\textsc{KL}}\left(\textrm{q}_{ij}(\cdot\mid u_{i},u_{j})\parallel\textrm{q}_{\mathrm{ref}}(\cdot\mid u_{i},u_{j})\right)\right]-\sum_{i=1}^{n}\mathbb{E}_{\textrm{q}_{\bm{u},i}(u_{i})}\left[\log\textrm{q}_{\bm{u},i}(u_{i})\right]\Bigg\},$
where $\textrm{q}_{\mathrm{ref}}(u_{ij}\mid u_{i},u_{j})\propto\exp\!\left(-\frac{\tau_{ij}}{2}\bigl[x_{ij}-\textrm{g}_{\mathrm{joint}}(u_{i},u_{j},u_{ij})\bigr]^{2}\right)$. Then, maximizing the ELBO over $\textrm{q}_{ij}$ for fixed $\textrm{q}_{\bm{u}}$ yields the stated objective. ∎
Proof of Eq.˜72.
Integrating over $\bm{z}_{n}\sim\mathcal{N}(0,\Sigma_{\theta})$ gives
$\bm{x}_{n}\mid\beta,\theta\sim\mathcal{N}\bigl(\bm{A}_{n}\beta,S_{\theta}\bigr).$
With $\pi(\beta)\equiv 1$, integrating over $\beta$ yields
$\log\mathrm{p}(\bm{x}_{n}\mid\theta)=-\tfrac{1}{2}\Bigl\{\log\det S_{\theta}+\log\bigl|\bm{A}_{n}^{\top}W_{\theta}\bm{A}_{n}\bigr|+(\bm{x}_{n}-\bm{A}_{n}\hat{\beta}_{\theta})^{\top}W_{\theta}(\bm{x}_{n}-\bm{A}_{n}\hat{\beta}_{\theta})\Bigr\}+C,$
where $C$ is constant in $\theta$. The objective $\ell(\bm{x}_{n};\theta)$ equals $\log\mathrm{p}(\bm{x}_{n}\mid\theta)$ up to an additive constant. ∎
Proof of Corollary˜1.
The discretized version of (59) is
$\sup_{\bm{w}\in\Delta_{K+1}^{\,n}}\ \sup_{\textrm{g}_{\mathrm{joint}}\in\mathcal{G}_{\mathrm{NN}}}\ \sum_{i=1}^{n}\sum_{k=0}^{K}w_{ik}\Bigg[$ $\widehat{\operatorname{LSE}}\!\left(x_{ii},\tau_{ii},u_{k},u_{k},\textrm{g}_{\mathrm{joint}}\right)+\sum_{j\neq i}\sum_{k_{2}=0}^{K}w_{jk_{2}}\,\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{k},u_{k_{2}},\textrm{g}_{\mathrm{joint}}\right)-\log w_{ik}\Bigg],$
where
$\widehat{\operatorname{LSE}}\!\left(x_{ij},\tau_{ij},u_{i},u_{j},\textrm{g}_{\mathrm{joint}}\right):=\log\sum_{k=0}^{K}\exp\!\left(-\frac{\tau_{ij}}{2}\left[x_{ij}-\textrm{g}_{\mathrm{joint}}(u_{i},u_{j},u_{k})\right]^{2}\right)-\log(K+1).$
Alternating maximization over $\bm{w}$ (softmax updates) and over $\textrm{g}_{\mathrm{joint}}$ (stochastic gradient step) yields the stated algorithm. ∎
Proof of Eq.˜75.
Conditional on $\beta$, the joint model for $\bm{z}_{S_{n}^{\prime}}$ and $\bm{x}_{S_{n}}$ is
$\bm{z}_{S_{n}^{\prime}}\sim\mathcal{N}\bigl(0,\Sigma_{\hat{\theta}}\bigr),\qquad\bm{x}_{S_{n}}\sim\mathcal{N}\bigl(\bm{A}_{S_{n}}\beta+\bm{z}_{S_{n}},D_{\tau}^{-1}\bigr).$
By Bayes’ rule,
$\mathrm{p}\bigl(\bm{z}_{S_{n}^{\prime}}\mid\bm{x}_{S_{n}},\beta\bigr)$ $\propto\exp\!\left(-\tfrac{1}{2}\bm{z}_{S_{n}^{\prime}}^{\top}\Sigma_{\hat{\theta}}^{-1}\bm{z}_{S_{n}^{\prime}}-\tfrac{1}{2}\bigl(\bm{x}_{S_{n}}-\bm{A}_{S_{n}}\beta-\bm{z}_{S_{n}}\bigr)^{\top}D_{\tau}\bigl(\bm{x}_{S_{n}}-\bm{A}_{S_{n}}\beta-\bm{z}_{S_{n}}\bigr)\right)$
$\propto\exp\!\left(-\tfrac{1}{2}\bm{z}_{S_{n}^{\prime}}^{\top}Q_{\hat{\theta}}\bm{z}_{S_{n}^{\prime}}+\bm{z}_{S_{n}^{\prime}}^{\top}L\right),$
where $Q_{\hat{\theta}}$ and $L$ are defined in the main text. Hence the posterior is Gaussian with mean $Q_{\hat{\theta}}^{-1}L$ and covariance $Q_{\hat{\theta}}^{-1}$. ∎
Appendix G Additional Empirical Results
G.1 Matrix Recovery Results
Table 3 reports the empirical Bayes matrix recovery results across all simulation settings described in §˜4.1.
Table 3: Comparison of matrix-recovery performance for NPMLE, EBMR–Sep, and EBMF across generating functions. We report the mean and standard error of R-MSE. The lowest R-MSE in each scenario is shown in bold.
n p $\bm{\tau}$ Linear Sine–log Sine–cos NPMLE EBMR-Sep EBMF NPMLE EBMR-Sep EBMF NPMLE EBMR-Sep EBMF 20 20 0.1 3.81 (0.48) 3.54 (0.38) 12.27 (2.23) 2.55 (0.47) 2.13 (0.29) 9.24 (0.00) 3.30 (0.43) 2.69 (0.41) 2.14 (0.00) 20 20 0.25 7.30 (0.22) 6.35 (0.35) 10.61 (0.40) 4.18 (0.23) 3.73 (0.21) 11.18 (1.38) 5.94 (0.27) 4.71 (0.30) 5.36 (0.00) 20 20 1 23.13 (0.57) 15.87 (0.72) 18.33 (0.50) 13.35 (0.39) 12.19 (0.66) 14.84 (0.48) 18.80 (0.34) 10.98 (1.41) 10.69 (1.29) 20 20 4 53.01 (1.28) 34.17 (0.65) 44.63 (0.78) 36.87 (0.96) 26.75 (0.80) 32.18 (0.75) 48.46 (0.91) 22.58 (7.24) 13.35 (0.71) 20 50 0.1 2.88 (0.08) 2.74 (0.17) 8.44 (0.30) 1.94 (0.07) 1.91 (0.16) 9.56 (0.55) 2.42 (0.09) 2.45 (0.22) 2.06 (0.00) 20 50 0.25 6.36 (0.08) 5.18 (0.18) 9.94 (0.31) 4.17 (0.08) 3.73 (0.25) 8.39 (0.23) 5.43 (0.11) 3.61 (0.29) 5.14 (0.00) 20 50 1 20.38 (0.21) 13.47 (1.29) 22.77 (0.38) 14.07 (0.16) 9.76 (0.65) 18.05 (0.42) 17.86 (0.21) 6.39 (0.42) 20.55 (0.00) 20 50 4 50.17 (0.67) 36.63 (6.97) 69.41 (0.23) 38.73 (0.51) 22.32 (0.47) 49.97 (0.22) 44.45 (0.51) 8.97 (0.55) 37.90 (1.30) 50 50 0.1 2.73 (0.06) 2.23 (0.04) 4.63 (0.14) 1.73 (0.06) 1.49 (0.05) 3.93 (0.15) 1.94 (0.05) 1.30 (0.10) 1.84 (0.00) 50 50 0.25 6.33 (0.09) 4.36 (0.12) 5.89 (0.15) 3.98 (0.08) 3.02 (0.07) 4.95 (0.14) 4.60 (0.09) 2.46 (0.37) 4.63 (0.22) 50 50 1 20.83 (0.14) 10.66 (0.16) 12.24 (0.18) 13.67 (0.11) 7.54 (0.16) 8.57 (0.18) 15.73 (0.09) 4.26 (1.56) 4.59 (0.17) 50 50 4 51.34 (0.50) 28.38 (0.30) 37.47 (0.30) 38.42 (0.31) 19.70 (0.18) 23.19 (0.28) 42.61 (0.37) 4.88 (0.27) 6.39 (0.24) n p $\bm{\tau}$ Tanh Reciprocal NPMLE EBMR-Sep EBMF NPMLE EBMR-Sep EBMF 20 20 0.1 1.45 (0.40) 1.06 (0.33) 7.57 (0.00) 1.36 (0.38) 1.02 (0.36) 1.84 (0.00) 20 20 0.25 1.25 (0.26) 0.79 (0.14) 11.35 (1.68) 1.14 (0.26) 1.36 (0.72) 4.61 (0.00) 20 20 1 2.81 (0.21) 2.64 (0.33) 10.16 (0.50) 1.84 (0.16) 1.92 (0.34) 12.46 (1.66) 20 20 4 8.99 (0.65) 7.14 (0.51) 14.49 (0.95) 5.26 (0.49) 4.54 (0.33) 12.04 (0.71) 20 50 0.1 0.59 (0.07) 0.56 (0.13) 7.47 (0.28) 0.57 (0.07) 0.58 (0.13) 1.75 (0.00) 20 50 0.25 0.94 (0.09) 0.54 (0.07) 5.75 (0.19) 0.78 (0.12) 0.41 (0.09) 4.38 (0.00) 20 50 1 2.13 (0.20) 1.75 (0.18) 8.07 (0.41) 1.62 (0.19) 1.11 (0.16) 8.16 (1.10) 20 50 4 5.60 (0.11) 4.88 (0.13) 10.65 (0.21) 3.18 (0.06) 3.14 (0.14) 9.01 (0.18) 50 50 0.1 0.41 (0.05) 0.32 (0.04) 3.91 (0.41) 0.29 (0.06) 0.25 (0.05) 2.14 (0.19) 50 50 0.25 0.77 (0.08) 0.60 (0.02) 3.91 (0.14) 0.50 (0.09) 0.34 (0.03) 4.21 (0.28) 50 50 1 2.31 (0.06) 2.01 (0.04) 4.85 (0.18) 1.23 (0.03) 1.01 (0.02) 4.13 (0.15) 50 50 4 8.07 (0.06) 5.78 (0.13) 7.87 (0.30) 4.07 (0.06) 3.43 (0.12) 5.23 (0.31)
G.2 Covariate-Assisted Empirical Bayes with Model Misspecification
To assess the robustness of Bayesian EB methods, we compare covariate-assisted empirical Bayes to competing methods under a heavy-tailed noise. For each $(n,p,T_{0})$ design, we generate independent noise $\varepsilon_{ij}$ from a scaled Student–$t$ distribution with $\nu=5$ degrees of freedom:
$\varepsilon_{ij}\sim c_{\nu}\,t_{\nu},\qquad c_{\nu}=\frac{\sigma}{\sqrt{\nu/(\nu-2)}},$
where $c_{\nu}$ is chosen to ensure $\operatorname{Var}(\varepsilon_{ij})=\sigma^{2}$. Thus, the noise variance matches the target variance but the tail probabilities are heavier than Gaussian.
Figure˜8 plots RMSE of $\hat{\bm{z}}_{n,p}$ versus $p$; points show medians across replicates and bands indicate interquartile ranges for each method.
Figure 8: CAEB under heavy-tailed noise. R-MSE of latent recovery versus dimension $p$ for three choices of g: linear, logistic, and nonlinear. Curves compare CAEB, EBMF, EBMR–sep, and NPMLE; points are medians and bands are IQR (25–75%). With noise following a $t_{5}$ distribution, CAEB attains the lowest error and continues to improve as the number of columns $p$ increases.
G.3 Additional Results for NYC Air Quality
Figure 9: Fitted spectral density for weekly $\mathrm{PM}_{2.5}$ in New York City, with $K=10$. The spectrum shows two dominant modes–one on the upper left corner and one on the lower right corner, which indicates two distinct spatio–temporal regimes. Based on the denoised series and site geography, one mode might correspond to the Manhattan sites (Herald Square, Chinatown, Lower East Side) and the other might correspond to northern Queens.
Figure 10: Posterior mean estimation of $\beta^{\star}$ improves with increasing $n$. Mean squared error (MSE) for estimating $\beta^{\star}$ using the posterior mean under the EB spatial regression, across varying sample size $n$ and precision $\tau$. We fit the spectral density using $K\in\{3,4,5\}$ Gaussian components; results are robust to moderate over-specification of $K$. Both axes use logarithmic scales. The points denote medians over 10 replicates, while bands show interquartile ranges (25–75%).
Figure 11: Posterior distributions of location effects in EB spatial regression for NYC $\mathrm{PM}_{2.5}$. Baseline is Herald Square (the mean is $7.654\,\mu\mathrm{g}/\mathrm{m}^{3}$). The effect of measuring at Lower East Side is statistically indistinguishable from the baseline. Queens North is lower by $1.357\,\mu\mathrm{g}/\mathrm{m}^{3}$ (posterior s.d. $0.224$), which provides strong evidence of better air quality there relative to Herald Square.